### v1

# Numerical methods for eigenvalues of singular polynomial eigenvalue problems

Abstract
Recently, three numerical methods for the computation of eigenvalues of singular matrix pencils, based on a rank-completing perturbation, a rank-projection, or an augmentation were developed. We show that all three approaches can be generalized to treat singular polynomial eigenvalue problems. The common denominator of all three approaches is a transformation of a singular into a regular matrix polynomial whose eigenvalues are a disjoint union of the eigenvalues of the singular polynomial, called true eigenvalues, and additional fake eigenvalues. The true eigenvalues can then be separated from the fake eigenvalues using information on the corresponding left and right eigenvectors. We illustrate the approaches on several interesting applications, including bivariate polynomial systems and ZGV points.
keywords:
Singular polynomial eigenvalue problem, perturbation, projection, augmentation, strong linearization, bivariate polynomial systems, ZGV points.
MSC:
65F15, 15A18, 15A22, 47A55, 65F22
1 Introduction
We consider the numerical computation of finite eigenvalues of the singular polynomial eigenvalue problem (singular PEP) associated with a matrix polynomial
$P(\lambda)=A_{0}+\lambda A_{1}+\cdots+\lambda^{d}A_{d}$ (1)
of degree $d\geq 2$, where $A_{0},\dots,A_{d}$, $A_{d}\neq 0$, are $m\times n$ (real or complex) matrices such that the polynomial $P(\lambda)$ is singular, which means that either $m=n$ and $\det\big{(}P(\lambda)\big{)}\equiv 0$, or $m\neq n$. Then $\lambda_{0}\in\mathbb{C}$ is a finite eigenvalue of $P$ if $\mathop{\rm rank}\nolimits\big{(}P(\lambda_{0})\big{)}<\mathop{\rm nrank}\nolimits(P)$, where
$\mathop{\rm nrank}\nolimits(P):=\max_{\zeta\in\mathbb{C}}\,\mathop{\rm rank}\nolimits\big{(}P(\zeta)\big{)}$
is the normal rank of the matrix polynomial $P$. (See Definition 2.3 for infinite eigenvalues.) The case $d=1$, where (1) is a singular matrix pencil, is covered in [10] and [11]. In this paper we show how we can extend the main ideas and tools to singular PEPs.
A standard approach to solve a PEP is to first linearize the matrix polynomial into a linear matrix pencil and then compute the eigenvalues (and possibly eigenvectors) from the obtained generalized eigenvalue problem. If we apply a strong linearization (see Section 2 for details) to a singular matrix polynomial $P$, we obtain a singular matrix pencil $L$ such that its finite and infinite eigenvalues agree with the eigenvalues of $P$. To compute the eigenvalues of $L$ we may, e.g., apply a numerical method from [10] or [11], where we perturb or project $L$ into a regular matrix pencil $\widetilde{L}$ such that all eigenvalues of $L$ are also eigenvalues of $\widetilde{L}$. Then the eigenvalues of $P$ can be extracted from the eigenvalues of $\widetilde{L}$ using orthogonality relations of the corresponding left and right eigenvectors.
In this paper we suggest an alternative approach. Instead of applying a linearization, we first perturb, project, or augment a singular matrix polynomial $P$ into a regular matrix polynomial $\widetilde{P}$ such that all eigenvalues of $P$ are also eigenvalues of $\widetilde{P}$. Then we compute the eigenvalues of $\widetilde{P}$ together with the left and right eigenvectors and extract the finite eigenvalues of $P$ from this set. To compute the eigenvalues of the regular matrix polynomial $\widetilde{P}$ we may apply any of the many numerical methods for such problems, including of course the use of linearizations and the solution of the corresponding generalized eigenvalue problem. In that case, one may wonder whether it is equivalent to first linearize $P$ into $L$ and then perturb (project or augment) $L$ into a regular pencil $\widetilde{L}$, or first perturb (project or augment) $P$ into a regular matrix polynomial $\widetilde{P}$ and then linearize $\widetilde{P}$ as a pencil $\widehat{L}$. We will show that these methods are not equivalent and that the second approach has the advantage of leading to a generalized eigenvalue problem of smaller size.
A different numerical method for singular quadratic eigenvalue problems (singular QEPs) has recently been proposed in [15]. In this method a singular QEP with $Q(\lambda)=\lambda^{2}M+\lambda C+K$ is perturbed into a regular problem with $\widetilde{Q}(\lambda)=\lambda^{2}\widetilde{M}+\lambda\widetilde{C}+\widetilde{K}$ by small random full rank perturbations. Next, the eigenvalues $\lambda_{i}$ of $\widetilde{Q}$ are computed together with the right and left eigenvectors and if the condition number of a computed eigenvalue is small enough, then the eigenvalue is identified as a finite eigenvalue of $Q$. The corresponding criterion is based on the notion of the $\delta$-weak condition number of an eigenvalue of a singular PEP, as introduced in [16].
The alternative approach proposed in this paper has an advantage that it leaves the eigenvalues intact and is therefore expected to return more accurate solutions. If we apply the projection variant, then another advantage are smaller matrices that require less computational work. A drawback of the new methods is that they rely on a correct determination of the normal rank. We give a more detailed comparison in Section 6.
The rest of the paper is organized as follows. Section 2 introduces the background theory on singular matrix polynomials. In Section 3 we generalize the main results from [10] and [11] for singular pencils to the setting of singular matrix polynomials. In Section 4 we then present the corresponding numerical methods that are directly applied to the given matrix polynomial and compare them in Section 5 to the approach that first applies a linearization. We illustrate the theoretical results and algorithms with numerical experiments in Section 6, where we also describe applications to bivariate polynomial systems and computation of ZGV points. Finally, we summarize some conclusions in Section 7.
2 Preliminaries
For the following theory on eigenvalues, nullspaces, and minimal indices of singular matrix polynomials we refer to, e.g., [3] or [4]. First, to define multiplicities of eigenvalues of a singular matrix polynomial, we recall the Smith form which is obtained under unimodular equivalence. We remind the reader that a square matrix polynomial is called unimodular if its determinant is a nonzero constant which is equivalent to saying that it is invertible when interpreted as a matrix over the field $\mathbb{C}(\lambda)$ of rational functions over $\mathbb{C}$.
Theorem 2.1 (Smith form)
Let $P(\lambda)$ be an $m\times n$ matrix polynomial of normal rank $r$. Then there exist unimodular matrix polynomials $E(\lambda)$ and $F(\lambda)$ of sizes $m\times m$ and $n\times n$, respectively, such that
$E(\lambda)\,P(\lambda)\,F(\lambda)=\mathop{\rm diag}\nolimits\big{(}d_{1}(\lambda),\dots,d_{r}(\lambda),0,\dots,0\big{)}=:D(\lambda),$ (2)
where $d_{1}(\lambda),\dots,d_{r}(\lambda)$ are monic polynomials such that $d_{j}(\lambda)$ is a divisor of $d_{j+1}(\lambda)$ for $j=1,\dots,r-1$. The $m\times n$ diagonal matrix polynomial $D(\lambda)$ is unique.
The nonzero diagonal elements $d_{1}(\lambda),\dots,d_{r}(\lambda)$ in the Smith form (2) are called the invariant polynomials of $P(\lambda)$.
Definition 2.2
Let $\lambda_{0}\in\mathbb{C}$ be a finite eigenvalue of a matrix polynomial $P(\lambda)$ of normal rank $r$. The invariant polynomials $d_{1}(\lambda),\dots,d_{r}(\lambda)$ of $P(\lambda)$ can then be uniquely factored as
$d_{i}(\lambda)=(\lambda-\lambda_{0})^{\alpha_{i}}\,p_{i}(\lambda),\ \textrm{where}\ \alpha_{i}\geq 0\ \ \textrm{and}\ \ p_{i}(\lambda_{0})\neq 0,$
for $i=1,\dots,r$. The exponents $0\leq\alpha_{1}\leq\cdots\leq\alpha_{r}$ are called the structural indices of $P(\lambda)$ at $\lambda_{0}$. The algebraic multiplicity of $\lambda_{0}$ is the sum $\alpha_{1}+\cdots+\alpha_{r}$ of the structural indices, while the geometric multiplicity is the number of nonzero structural indices.
Definition 2.3
Let $P(\lambda)$ be a matrix polynomial (1) of degree $d$. Then
${\rm rev}\,P(\lambda):=\lambda^{j}\,P(1/\lambda)=A_{d}+\lambda A_{d-1}+\cdots+\lambda^{d}A_{0}$
is the reversal of $P(\lambda)$. We say that $\lambda_{0}=\infty$ is an eigenvalue of $P(\lambda)$ if 00 is an eigenvalue of ${\rm rev}\,P(\lambda)$. The structural indices of $P(\lambda)$ at $\lambda_{0}=\infty$ and the algebraic and geometric multiplicity of eigenvalue $\lambda_{0}=\infty$ are then defined as structural indices, and the algebraic and geometric multiplicity of ${\rm rev}\,P(\lambda)$ at 00, respectively.
Definition 2.4
The right and left nullspaces of an $m\times n$ matrix polynomial $P(\lambda)$ are the vector spaces
${\cal N}_{r}(P)$ =$=\{\,x(\lambda)\in\mathbb{C}(\lambda)^{n}:P(\lambda)\,x(\lambda)\equiv 0\,\},$
${\cal N}_{l}(P)$ =$=\{\,y(\lambda)\in\mathbb{C}(\lambda)^{m}:y(\lambda)^{*}P\,(\lambda)\equiv 0\,\}$
of rational vectors $x(\lambda)$ and $y(\lambda)$ annihilated by $P(\lambda)$.
Clearly, the following identity holds:
$\mathop{\rm nrank}\nolimits(P)=n-\dim{\cal N}_{r}(P)=m-\dim{\cal N}_{l}(P).$
For square matrix polynomials, i.e., when $m=n$, it then follows that $\dim{\cal N}_{r}(P)=\dim{\cal N}_{l}(P)$.
A basis of a subspace ${\cal V}$ of $\mathbb{C}(\lambda)^{n}$ is called a polynomial basis [6]. A polynomial basis is called minimal if the sum of the degrees of its polynomials is minimal along all polynomial bases of ${\cal V}$.
Definition 2.5
Let $P(\lambda)$ be an $m\times n$ singular matrix polynomial of normal rank $r$. Let $\{x_{1}(\lambda),\dots,x_{n-r}(\lambda)\}$ and $\{y_{1}(\lambda),\dots,y_{m-r}(\lambda)\}$ be minimal bases of, respectively, the right and left nullspaces of $P(\lambda)$, ordered so that $\varepsilon_{1}\leq\cdots\leq\varepsilon_{n-r}$ and $\eta_{1}\leq\cdots\leq\eta_{m-r}$, where $\varepsilon_{i}=\deg(x_{i})$ for $i=1,\dots,n-r$ and $\eta_{i}=\deg(y_{i})$ for $i=1,\dots,m-r$. Then $\varepsilon_{1},\dots,\varepsilon_{n-r}$ and $\eta_{1},\dots,\eta_{m-r}$ are, respectively, the right and left minimal indices of $P(\lambda)$.
The minimal bases are not unique, but the minimal indices are. If $\{x_{1}(\lambda),\dots,x_{n-r}(\lambda)\}$ and $\{y_{1}(\lambda),\dots,y_{m-r}(\lambda)\}$ are minimal bases for the right and left nullspaces of $P(\lambda)$, then the right and left singular spaces at $\mu\in\mathbb{C}$ are defined as
$\mathop{\rm Ker}\nolimits_{\mu}\big{(}P(\lambda)\big{)}$ $:={\rm span}\big{(}x_{1}(\mu),\dots,x_{n-r}(\mu)\big{)},$
$\mathop{\rm Ker}\nolimits_{\mu}\big{(}P(\lambda)^{*}\big{)}$ $:={\rm span}\big{(}y_{1}(\mu),\dots,y_{m-r}(\mu)\big{)}.$
The right and left singular spaces at $\mu$ do not depend on the choice of the minimal bases [5, Lemma 2.8].
The index sum theorem from [4, Lemma 6.3] will play an important role in the next section. The result can be used to compute the number of eigenvalues of a matrix polynomial if the minimal indices are known.
Theorem 2.6 (Index sum theorem)
Let $P(\lambda)$ be an $m\times n$ singular matrix polynomial of degree $d$ and normal rank $r$ with right minimal indices $\varepsilon_{1}\leq\cdots\leq\varepsilon_{n-r}$ and left minimal indices $\eta_{1}\leq\cdots\leq\eta_{m-r}$. Then
$d\,r=\delta_{\rm fin}(P)+\delta_{\infty}(P)+\sum_{i=1}^{n-r}\varepsilon_{i}+\sum_{i=1}^{m-r}\eta_{i},$ (3)
where $\delta_{\rm fin}(P)$ is the sum of the algebraic multiplicities of all finite eigenvalues of $P$ and $\delta_{\rm\infty}(P)$ is the algebraic multiplicity of eigenvalue $\infty$.
If $\lambda_{0}\in\mathbb{C}$ is an eigenvalue of an $m\times n$ singular matrix polynomial $P(\lambda)$ of degree $d$ and normal rank $r$, then $\dim\big{(}\mathop{\rm Ker}\nolimits(P(\lambda_{0}))\big{)}>n-r$ and as a representative of a right eigenvector we can take each $x_{0}\in[x]\in\mathop{\rm Ker}\nolimits\big{(}P(\lambda_{0})\big{)}\,/\,\mathop{\rm Ker}\nolimits_{\lambda_{0}}\big{(}P(\lambda)\big{)}$, see, e.g., [5] or [16]. In a similar way each $y_{0}\in[y]\in\mathop{\rm Ker}\nolimits\big{(}(P(\lambda_{0})^{*})\big{)}\,/\,\mathop{\rm Ker}\nolimits_{\lambda_{0}}\big{(}P(\lambda)^{*}\big{)}$ is a representative of a left eigenvector. If $\lambda_{0}$ is geometrically simple, i.e., $\dim\big{(}\mathop{\rm Ker}\nolimits(P(\lambda_{0}))\big{)}=n-r+1$, then we can uniquely (up to a scalar) define the right eigenvector and left eigenvector by the additional requirement that $x$ and $y$ are respectively orthogonal to $\mathop{\rm Ker}\nolimits_{\lambda_{0}}\big{(}P(\lambda)\big{)}$ and $\mathop{\rm Ker}\nolimits_{\lambda_{0}}\big{(}P(\lambda)^{*}\big{)}$.
The definition of the condition number for a finite eigenvalue is based on a restriction of the expansion result from [2, Thm. 2 and Eq. (18)] to the case of a simple eigenvalue; see also [16, Thm. 3.2].
Theorem 2.7
Let $\lambda_{0}$ be a finite simple eigenvalue of an $n\times n$ matrix polynomial $P(\lambda)$ of degree $d$ that has normal rank $r$. Let $X=[X_{1}\ \,x]$ be an $n\times(n-r+1)$ matrix with orthonormal columns such that the columns of $X_{1}$ form a basis for $\mathop{\rm Ker}\nolimits_{\lambda_{0}}\big{(}P(\lambda)\big{)}$ and the columns of $X$ form a basis for $\mathop{\rm Ker}\nolimits\big{(}P(\lambda_{0})\big{)}$, and let $Y=[Y_{1}\ \,y]$ be an $n\times(n-r+1)$ matrix with orthonormal columns such that the columns of $Y_{1}$ form a basis for $\mathop{\rm Ker}\nolimits_{\lambda_{0}}\big{(}P(\lambda)^{*}\big{)}$ and the columns of $Y$ form a basis for ${\mathop{\rm Ker}\nolimits}\big{(}P(\lambda_{0})^{*}\big{)}$. If $E(\lambda)$ is an $n\times n$ matrix polynomial of degree $d$ such that $Y_{1}^{*}E(\lambda_{0})\,X_{1}$ is nonsingular, then, for sufficiently small $\varepsilon>0$, there exists an eigenvalue $\lambda_{0}(\varepsilon)$ of the perturbed pencil $P(\lambda)+\varepsilon\,E(\lambda)$ such that
$\lambda_{0}(\varepsilon)=\lambda_{0}-\frac{\det(Y^{*}E(\lambda_{0})\,X)}{(y^{*}P^{\prime}(\lambda_{0})\,x)\cdot\det(Y_{1}^{*}E(\lambda_{0})\,X_{1})}\,\varepsilon+{\cal O}(\varepsilon^{2}).$ (4)
In the regular case, when $n=r$, the above result simplifies into a well-known expression for the eigenvalue expansion of a simple eigenvalue of a regular matrix polynomial [20], namely
$\lambda_{0}(\varepsilon)=\lambda_{0}-\frac{y^{*}E(\lambda_{0})\,x}{y^{*}P^{\prime}(\lambda_{0})\,x}\,\varepsilon+{\cal O}(\varepsilon^{2}).$ (5)
Based on Theorem 2.7 the condition number of a simple eigenvalue can be generalized to singular matrix polynomials.
Definition 2.8
Let $\lambda_{0}$ be a finite simple eigenvalue of an $n\times n$ singular matrix polynomial $P(\lambda)$ of degree $d$ that has normal rank $r$. Let $X=[X_{1}\ \,x]$ be an $n\times(n-r+1)$ matrix with orthonormal columns such that columns of $X_{1}$ form a basis for $\mathop{\rm Ker}\nolimits_{\lambda_{0}}\big{(}P(\lambda)\big{)}$ and columns of $X$ form a basis for $\mathop{\rm Ker}\nolimits\big{(}P(\lambda_{0})\big{)}$, and let $Y=[Y_{1}\ \,y]$ be an $n\times(n-r+1)$ matrix with orthonormal columns such that columns of $Y_{1}$ form a basis for $\mathop{\rm Ker}\nolimits_{\lambda_{0}}\big{(}P(\lambda)^{*}\big{)}$ and columns of $Y$ form a basis for ${\mathop{\rm Ker}\nolimits}\big{(}P(\lambda_{0})^{*}\big{)}$. Then we define
$\gamma(\lambda_{0})=|y^{*}P^{\prime}(\lambda_{0})\,x|\cdot\big{(}1+|\lambda_{0}|^{2}+\cdots+|\lambda_{0}|^{2d}\big{)}^{-1/2}$ (6)
and consider $\kappa(\lambda_{0})=\gamma(\lambda_{0})^{-1}$ as the condition number of $\lambda_{0}$.
In the regular case $\kappa(\lambda_{0})$ is indeed the condition number of $\lambda_{0}$. Namely, if we perturb $P(\lambda)=A_{0}+\lambda A_{1}+\cdots+\lambda^{d}A_{d}$ into $\widetilde{P}(\lambda)=P(\lambda)+\varepsilon E(\lambda)$, where $E(\lambda)=E_{0}+\lambda E_{1}+\cdots+\lambda^{d}E_{d}$ with $\|E\|_{*}:=\left(\|E_{0}\|^{2}+\cdots+\|E_{d}\|^{2}\right)^{1/2}\leq 1$, and $\lambda_{0}(\varepsilon)$ is the corresponding perturbed eigenvalue of $\widetilde{P}$, then
$|\lambda_{0}(\varepsilon)-\lambda_{0}|\leq\kappa(\lambda_{0})\,\varepsilon+{\cal O}(\varepsilon^{2})\,;$
see, e.g., [20] or [15].
On the other hand, if $\lambda_{0}\in\mathbb{C}$ is an eigenvalue of a singular matrix polynomial $P$, then arbitrarily small perturbations can move $\lambda_{0}$ to an arbitrary complex number as $\det(Y_{1}^{*}E(\lambda_{0})X_{1})$ in (4) can be arbitrarily close to 00. However, as explained by Lotz and Noferini [16], such behavior occurs very rarely in practice. If we exclude a set of small measure $\delta>0$ from all possible perturbations such that $\|E\|_{*}\leq 1$, then there exists a constant $C\geq 1$, which only depends on $m$, $n$, $d$, and $r$, such that
$|\lambda_{0}(\varepsilon)-\lambda_{0}|\leq C\,\delta^{-1}\,\kappa(\lambda_{0})\,\varepsilon+{\cal O}(\varepsilon^{2}),$
and $\kappa(\lambda_{0})$ is the quantity that describes how ill conditioned an eigenvalue is in this weak context. This is the reason why we will later use an approximation of (6) as a criterion whether or not a computed eigenvalue is a finite eigenvalue of a singular matrix polynomial $P$.
The following definition of a linearization of matrix polynomials can be found, e.g., in [8], while the concept of strong linearizations has been introduced in [7].
Definition 2.9
A matrix pencil $L(\lambda)=\lambda X+Y$ with $X,Y\in\mathbb{C}^{nd\times nd}$ is a linearization of an $n\times n$ matrix polynomial $P(\lambda)$ of degree $d$ if there exist two unimodular $nd\times nd$ matrix polynomials $E(\lambda)$ and $F(\lambda)$ such that
$E(\lambda)\,L(\lambda)\,F(\lambda)=\left[\begin{matrix}P(\lambda)&0\cr 0&I_{(d-1)n}\end{matrix}\right].$
A linearization $L(\lambda)$ is called a strong linearization if the reversal polynomial ${\rm rev}\,L(\lambda)$ is also a linearization of the reversal polynomial ${\rm rev}\,P(\lambda)$.
It follows immediately from the definition that a matrix polynomial and its linearization have the same Smith form, so the finite eigenvalues and the corresponding structural indices of the matrix polynomial are exactly the finite eigenvalues and corresponding structural indices of its linearization. If the linearization is strong, then this also holds for the eigenvalue $\infty$.
We say that a set $\mathcal{A}\subseteq\mathbb{C}^{n}$ is algebraic if it is the set of common zeros of finitely many polynomials. A set $\Omega\subseteq\mathbb{C}^{n}$ is called generic if its complement is contained in an algebraic set that is not the full space $\mathbb{C}^{n}$. A set $\Omega\subseteq\mathbb{C}^{n\times k}$ is called generic if it can be canonically identified with a generic subset of $\mathbb{C}^{nk}$. We use the expression that a property holds generically with respect to the entries of $U\in\mathbb{C}^{n\times k}$ if there exists a generic set $\Omega\subseteq\mathbb{C}^{n\times k}$ such that the property holds for all $U\in\Omega$.
3 Main results
The main results are generalizations of similar results from [10] and [11]. The corresponding proofs of the results in the pencil case, however, rely on the existence of the Kronecker canonical form for matrix pencils. Since an analogous canonical form under strict equivalence is not available for matrix polynomials, we need different arguments to extend the results to the polynomial case.
Lemma 3.1
Let $P(\lambda)$ be an $n\times n$ singular matrix polynomial of degree $d$ and normal rank $n-k$, and let $S_{R}(\lambda)=\left[x_{1}(\lambda)\ \dots\ x_{k}(\lambda)\right]$ and $S_{L}(\lambda)=\left[y_{1}(\lambda)\ \dots\ y_{k}(\lambda)\right]$ be $n\times k$ matrix polynomials, such that the columns form minimal bases of, respectively, the right and left nullspaces of $P(\lambda)$ with left minimal indices $\eta_{1},\dots,\eta_{k}$ and right minimal indices $\varepsilon_{1},\dots,\varepsilon_{k}$. Furthermore, let $U,V\in\mathbb{C}^{n\times k}$ have full column rank, $N:=\eta_{1}+\cdots+\eta_{k}$, $M:=\varepsilon_{1}+\cdots+\varepsilon_{k}$, and let $\gamma_{1},\dots,\gamma_{dk}\in\mathbb{C}$ be given values that are distinct from the eigenvalues of $P(\lambda)$. Then, generically with respect to the entries of $U$ and $V^{*}$, the following statements hold:
  1. 1)
    The polynomial $\det\big{(}V^{*}S_{R}(\lambda)\big{)}$ has exactly $M$ simple roots $\alpha_{1},\dots,\alpha_{M}$ that are different from the eigenvalues of $P(\lambda)$ and different from $\gamma_{1},\dots,\gamma_{dk}$. For each $\alpha_{i}$ there exists a nonzero vector $z_{i}$ such that $P(\alpha_{i})\,z_{i}=0$ and $V^{*}z_{i}=0$.
  2. 2)
    The polynomial $\det\big{(}U^{*}S_{L}(\lambda)\big{)}$ has exactly $N$ simple roots $\beta_{1},\dots,\beta_{N}$ that are different from the eigenvalues of $P(\lambda)$, from the values $\gamma_{1}$, …, $\gamma_{dk}$, and also from $\alpha_{1}$, …, $\alpha_{M}$. For each $\beta_{i}$ there exists a nonzero vector $w_{i}$ such that $w_{i}^{*}P(\beta_{i})=0$ and $w_{i}^{*}U=0$.
  3. 3)
    For any $dk$ nonzero vectors $t_{1},\dots,t_{dk}\in\mathbb{C}^{k}$ there exist nonzero vectors $s_{1},\dots,s_{dk}\in\mathbb{C}^{n}$ with $P(\gamma_{i})\,s_{i}=0$ and $t_{i}=V^{*}s_{i}$ for $i=1,\dots,dk$.
Proof 1
1) The $k\times k$ matrix $G(\alpha):=V^{*}S_{R}(\alpha)$ has elements $g_{ij}(\alpha)=v_{i}^{*}x_{j}(\alpha)$ that are polynomials in $\alpha$ which, generically with respect to the entries of $v_{i}^{*}$, have degree $\varepsilon_{j}$ for $i,j=1,\dots,k$. It follows that $\det G(\alpha)$ is a polynomial in $\alpha$ which generically with respect to the entries of $V^{*}$ is of degree $M$ (cf. the main theorem in [6]). Thus $\det G(\alpha)=0$ has $M$ roots $\alpha_{1},\dots,\alpha_{M}$ (counted with multiplicities).
For each root $\alpha_{i}$ there exists a nonzero vector $s_{i}\in\mathbb{C}^{k}$ such that $G(\alpha_{i})\,s_{i}=0$. We know (see, e.g., [21, Thm. 2.2]) that $S_{R}(\alpha)$ is of full column rank for all $\alpha$. So, if we take $z_{i}=S_{R}(\alpha_{i})\,s_{i}$, then $z_{i}$ is nonzero and $V^{*}z_{i}=V^{*}S_{R}(\alpha_{i})\,s_{i}=G(\alpha_{i})\,s_{i}=0$. On the other hand, since $z_{i}\in\mathop{\rm Ker}\nolimits_{\alpha_{i}}\big{(}P(\lambda)\big{)}$, it also holds that $P(\alpha_{i})\,z_{i}=0$.
For a fixed $\mu\in\mathbb{C}$ we can consider $\det G(\mu)$ as a polynomial in the entries of $V^{*}=[v_{1}\ \,\dots\ \,v_{k}]^{*}$. For the particular choice $V=S_{R}(\mu)$ we obtain that $G(\mu)$ is nonsingular, which shows that $\det G(\mu)$ is a nonzero polynomial in the entries of $V^{*}$. It follows that $\det G(\mu)\neq 0$ generically with respect to the entries of $V^{*}$, and consequently the fixed value $\mu$ will generically not be among the roots of $\det G(\alpha)$ as a polynomial in $\alpha$. Since the intersection of finitely many generic sets is still generic, it follows that we can exclude finitely many values from the roots $\alpha_{1},\dots,\alpha_{M}$ of $G(\alpha)$. This shows that generically with respect to the entries of $V^{*}$, the values $\alpha_{1},\dots,\alpha_{M}$ are different from the eigenvalues of $P(\lambda)$ and also from the given values $\gamma_{1},\dots,\gamma_{dk}$.
It remains to show that the roots $\alpha_{1},\dots,\alpha_{M}$ of $\det G(\alpha)$ are generically simple. This is the case if the discriminant of $\det G(\alpha)$, which is a polynomial in the entries of $V^{*}$, is nonzero. Thus, to show that the roots of $\det G(\alpha)$ are generically simple, it remains to show that this discriminant is not the zero polynomial and to this end it is sufficient to find one particular example for $V$. The existence of such a $V$ for which the roots of $\det G(\alpha)$ are simple is guaranteed by Theorem A.3 in the appendix.
2) Similarly as in 1) we now consider the left null space of $P(\lambda)$ and show the existence of $\beta_{1},\dots,\beta_{N}$ and the corresponding nonzero vectors $w_{1},\dots,w_{N}$, where now the statements are generic with respect to the entries of $U$. In particular, by interpreting $V$ as already fixed, this shows that generically with respect to the entries of $U$, the values $\beta_{1},\dots,\beta_{N}$ are not only different from the eigenvalues of $P(\lambda)$ and $\gamma_{1},\dots,\gamma_{dk}$, but also from the values $\alpha_{1},\dots,\alpha_{M}$ constructed in 1)1)1 ).
3) With the same notation as in 1) we now aim to solve the equations $G(\gamma_{i})\,c_{i}=t_{i}$ for $i=1,\dots,dk$. Since $\gamma_{i}$ is different from the values $\alpha_{1},\dots,\alpha_{M}$, we have $\det G(\gamma_{i})\neq 0$ and hence $G(\gamma_{i})\,c_{i}=t_{i}$ is uniquely solvable for nonzero $c_{i}$ for $i=1,\dots,dk$. The vectors that we are looking for are $s_{i}=S_{R}(\gamma_{i})\,c_{i}$ for $i=1,\dots,dk$. $\qed$
Theorem 3.2
Let $P(\lambda)=A_{0}+\lambda A_{1}+\cdots+\lambda^{d}A_{d}$ be an $n\times n$ singular matrix polynomial of degree $d$ and normal rank $n-k$ with right minimal indices $\varepsilon_{1}\leq\cdots\leq\varepsilon_{k}$ and left minimal indices $\eta_{1}\leq\cdots\leq\eta_{k}$ such that all its eigenvalues are semisimple, and let $M=\varepsilon_{1}+\cdots+\varepsilon_{k}$ and $N=\eta_{1}+\cdots+\eta_{k}$. Furthermore, let $B_{0},\dots,B_{d}\in\mathbb{C}^{k\times k}$ be such that the matrix polynomial $Q(\lambda):=B_{0}+\lambda B_{1}+\cdots+\lambda^{d}B_{d}$ is regular and its eigenvalues are simple and distinct from the eigenvalues of $P(\lambda)$. Then there exists a generic set $\Omega\subseteq\mathbb{C}^{n\times k}\times\mathbb{C}^{k\times n}$ such that for all $(U,V^{*})\in\Omega$ and all $\tau>0$ the matrix polynomial
$\widetilde{P}(\lambda)=P(\lambda)+\tau\,U\,Q(\lambda)\,V^{*},$ (7)
is regular and its $nd$ eigenvalues are independent of $\tau$ and can be classified into the following four groups:
  1. (a)
    (True eigenvalues) If $\lambda_{0}$ is an eigenvalue of $P(\lambda)$, then $\lambda_{0}$ is also an eigenvalue of $\widetilde{P}(\lambda)$. If $x$ and $y$ are corresponding right and left eigenvectors then $V^{*}x=0$ and $U^{*}y=0$. There are $(n-k)d-M-N$ such eigenvalues.
  2. (b)
    (Right random eigenvalues) There are $M$ simple eigenvalues, which are all different from the eigenvalues in (a) and those of $Q(\lambda)$, such that $V^{*}x=0$ and $U^{*}y\neq 0$, where $x$ and $y$ are corresponding right and left eigenvectors.
  3. (c)
    (Left random eigenvalues) There are $N$ simple eigenvalues, which are all different from the eigenvalues in (a) and (b) and those of $Q(\lambda)$, such that $V^{*}x\neq 0$ and $U^{*}y=0$, where $x$ and $y$ are corresponding right and left eigenvectors.
  4. (d)
    (Prescribed eigenvalues) If $\lambda_{0}$ is an eigenvalue of $Q(\lambda)$ and if $x$ and $y$ are corresponding right and left eigenvector, then $V^{*}x\neq 0$ and $U^{*}y\neq 0$. There are $kd$ such eigenvalues.
Remark 3.3
We use the attribute “right” or “left” in (a) and (b) to indicate that the corresponding random eigenvalues can be related to the right or left minimal indices of the polynomial $P(\lambda)$, respectively.
Proof 2
First, we note that $U$ and $V$ generically have full column rank, so in the following we assume that this is the case. Suppose that $\mu_{0}$ is neither an eigenvalue of $P(\lambda)$ nor $Q(\lambda)$, which means that $\mathop{\rm rank}\nolimits\big{(}P(\mu_{0})\big{)}=n-k$. Then we have $\det\big{(}\widetilde{P}(\mu_{0})\big{)}\neq 0$ for all $\tau>0$ if and only if the range of $UQ(\mu_{0})V^{*}$ has a trivial intersection with the kernel of $P(\mu)$. By the dimension formula, it is easy to construct matrices $U,V\in\mathbb{C}^{n\times k}$ such that this is the case. Since $\det\big{(}\widetilde{P}(\mu_{0})\big{)}$ is a polynomial in the entries of $U$ and $V^{*}$, this polynomial is not identical to zero and it follows that generically $\mathop{\rm nrank}\nolimits(\widetilde{P})=n$ independent of $\tau$. In such case $\widetilde{P}(\lambda)$ has $nd$ eigenvalues (some of them can be infinite if $A_{d}+\tau\,UB_{d}V^{*}$ is singular).
(a) Let $\lambda_{0}$ be a finite eigenvalue of $P(\lambda)$ of multiplicity $r\geq 1$. Then $\dim\big{(}\mathop{\rm Ker}\nolimits(P(\lambda_{0}))\big{)}=k+r$ and there exist $r$ linearly independent vectors $x_{1},\dots,x_{r}\in\mathop{\rm Ker}\nolimits(P(\lambda_{0}))$ such that $V^{*}x_{i}=0$ for $i=1,\dots,r$. It follows that $\widetilde{P}(\lambda_{0})\,x_{i}=0$ for $i=1,\dots,r$ and hence $\lambda_{0}$ is an eigenvalue of $\widetilde{P}(\lambda)$ of multiplicity at least $r$ independent of $\tau>0$. In a similar way we see that there exist $r$ linearly independent vectors $y_{1},\dots,y_{r}\in\mathop{\rm Ker}\nolimits(P(\lambda_{0})^{*})$ such that $U^{*}y_{i}=0$ and $y_{i}^{*}P(\lambda_{0})=y_{i}^{*}\widetilde{P}(\lambda_{0})=0$.
To show that (a) holds for $\lambda_{0}=\infty$ as well, we apply the already proved part to the reversal ${\rm rev}\,P(\lambda)$.
(b) In the following let $\gamma_{1},\dots,\gamma_{dk}$ denote the eigenvalues of $Q(\lambda)$. Let $S_{R}(\lambda)$ be an $n\times k$ matrix polynomial such that its columns form a minimal basis of the right nullspace ${\cal N}_{r}$ of $P(\lambda)$. By Lemma 3.1, $\det(V^{*}S_{R}(\alpha))$ generically is a polynomial of degree $M$ with simple roots $\alpha_{1},\dots,\alpha_{M}$ which are all different from the eigenvalues of $P(\lambda)$ and from $\gamma_{1},\dots,\gamma_{dk}$. If $\alpha_{0}$ is such an root, then by Lemma 3.1 there exists a nonzero vector $z$ such that $P(\alpha_{0})\,z=0$ and $V^{*}z=0$. It follows that $\widetilde{P}(\alpha_{0})\,z=0$, thus $\alpha_{0}$ is an eigenvalue of $\widetilde{P}$ and $z$ is a right eigenvector, independent of $\tau>0$.
Considering now $\tau$ as a variable, it follows that the pencil
$G+\tau H:=P(\alpha_{0})+\tau\,U\,Q(\alpha_{0})\,V^{*}$ (8)
is singular. Clearly, we have $\mathop{\rm rank}\nolimits(G)=n-k$ and $\mathop{\rm rank}\nolimits(H)=k$, because $\alpha_{0}$ differs from all eigenvalues of $P(\lambda)$ and $Q(\lambda)$ and because $U$ and $V$ have full column rank.
Suppose that ${\rm nrank}(G,H)=n-j$ for $j\geq 1$, which means that (8) has $j$ right and $j$ left minimal indices. We know from $Gz=0$ and $Hz=0$ that at least one right minimal index is equal to zero. The remaining $j-1$ right minimal indices are all larger than zero, because otherwise there would exist a nonzero $y\in\mathop{\rm Ker}\nolimits(G)\cap\mathop{\rm Ker}\nolimits(H)$ linearly independent of $z$. But then $Gy=0$ implies that $y$ is in the range of $S_{R}(\alpha_{0})$, i.e., there exists $\widetilde{y}\in\mathbb{C}^{k}$ such that $y=S_{R}(\alpha_{0})\widetilde{y}$. Thus, since $Hy=0$, we have that $V^{*}S_{R}(\alpha_{0})\widetilde{y}=0$, because $U$ has full column rank and $Q(\alpha_{0})$ is nonsingular. But this implies $\alpha_{0}$ would be a multiple root of $\det V^{*}S_{R}(\alpha)$, which is not possible due to Lemma 3.1.
Now suppose that (8) has a left minimal index zero. Then there exists a nonzero vector $w$ such that $w^{*}G=0$ and $w^{*}H=0$, which implies $w^{*}U=0$ because $V$ has full rank and $Q(\alpha_{0})$ is nonsingular. But then $\alpha_{0}$ is equal to one of the values $\beta_{1},\dots,\beta_{N}$ from Lemma 3.1 which is a contradiction to Lemma 3.1. Thus, all left minimal indices of (8) are larger than or equal to one.
Now $\mathop{\rm rank}\nolimits(H)=k$ implies that $\tau_{0}=\infty$ is an eigenvalue of the pencil (8) of geometric multiplicity $n-k-j$ and thus algebraic multiplicity at least $n-k-j$. Similarly, $\tau_{0}=0$ is an eigenvalue of algebraic multiplicity at least $k-j$. Therefore, the sum of the multiplicities of the finite and infinite eigenvalues of (8) is at least $n-2j$. On the other hand, the normal rank of (8) is $n-j$ and the pencil has $j$ left minimal indices that are all greater or equal to one. It follows from (3) of the index sum theorem that the only possible option is that $j=1$. Therefore, (8) has one left minimal index one, one right minimal index zero, and 00 and $\infty$ are the only eigenvalues of (8).
It follows that there exist linearly independent vectors $w$ and $z$ such that
$w^{*}G=0,\quad z^{*}G=w^{*}H\neq 0,\quad z^{*}H=0,$ (9)
which implies $U^{*}w\neq 0$. Up to scaling, the left eigenvector $y$ of (7) associated with $\alpha_{0}$ then has the form $y(\tau)=w+\tau z$, is a linear function of $\tau$, and satisfies $U^{*}y(\tau)\neq 0$.
(c) We show this in a similar way as (b).
(d) Let $\gamma_{i}$ be an eigenvalue of $Q(\lambda)$ with the right eigenvector $t_{i}\in\mathbb{C}^{k}$. It follows from Lemma 3.1 that there exists a nonzero vector $x_{i}$ such that $P(\gamma_{i})\,x_{i}=0$ and $t_{i}=V^{*}x_{i}$, therefore $\widetilde{P}(\gamma_{i})\,x_{i}=0$ and $V^{*}x_{i}\neq 0$ for $i=1,\dots,dk$ independent of $\tau>0$. In a similar way we can find a left eigenvector $y_{i}$ such that $y_{i}^{*}P(\gamma_{i})=0$ and $U^{*}y_{i}\neq 0$, again independent of $\tau>0$. Noting that due to the simplicity of the eigenvalues of $\gamma_{i}$ left and right eigenvectors are unique up to multiplication with nonzero scalars concludes the proof of (d).
By counting the number of eigenvalues of $\widetilde{P}(\lambda)$, we can see from Theorem 2.6 (index sum theorem) that there are no other eigenvalues than those from (a), (b), (c), and (d), and in addition, that the algebraic multiplicity of each of the eigenvalues in (a) are equal for the pencils $P(\lambda)$ and $\widetilde{P}(\lambda)$$\qed$
We highlight that the right and left random eigenvalues from Theorem 3.2 depend only on the choices of $U$ and $V$, respectively. In particular, they do not depend on the choices of the matrix polynomial $Q(\lambda)$ and of $\tau>0$.
Remark 3.4
The corresponding result in the pencil case [10] has no restrictions on the multiplicity of the eigenvalues of the singular pencil, because the corresponding proof is based on the Kronecker canonical form of matrix pencils which allows the separation of the regular part from the singular part of the matrix pencil. For matrix polynomials there unfortunately is no straightforward extension, and the alternative proof strategy in Theorem 3.2 needs the additional assumption that the eigenvalues of the singular pencil are semisimple. However, numerical experiments suggest that the result also holds for singular matrix polynomials that have eigenvalues that are not semisimple.
Let us consider a given singular matrix polynomial $P(\lambda)$ and matrices $U$, $V$, $B_{0},\dots,B_{k}$ as in Theorem 3.2. If $U_{\perp}$ and $V_{\perp}$ are matrices whose columns form bases for the orthogonal complement of the ranges of $U$ and $V$, respectively, then we can build a strictly equivalent matrix polynomial
$\widehat{P}(\lambda):=[U\ \,U_{\perp}]^{*}\,P(\lambda)\,[V\ \,V_{\perp}]=\left[\!\!\begin{array}[]{cc}U^{*}P(\lambda)\,V&U^{*}P(\lambda)\,V_{\perp}\\[2.84526pt] U_{\perp}^{*}P(\lambda)\,V&U_{\perp}^{*}P(\lambda)\,V_{\perp}\end{array}\!\!\right].$ (10)
In comparison with (10), the perturbed matrix polynomial (7) only differs from the original pencil by an extra term $\tau\,Q(\lambda)$ in the $(1,1)$-block. The following generalization of [11, Prop. 4.1] shows that under mild additional assumptions it is possible to extract the true eigenvalues of the matrix polynomial $P(\lambda)$ from the projected matrix polynomial $U_{\perp}^{*}P(\lambda)\,V_{\perp}$ of size $(n-k)\times(n-k)$. $\qed$
Proposition 3.5
Let $P(\lambda)$ be a complex $n\times n$ matrix polynomial of degree $d$ and normal rank $n-k$ such that all its eigenvalues are semisimple. Furthermore, let $U$, $V$, $B_{0},\dots,B_{k}$ satisfy the hypotheses of Theorem 3.2. If the $(n-k)\times(n-k)$ matrix polynomial $P_{22}(\lambda):=U_{\perp}^{*}P(\lambda)\,V_{\perp}$ from (10) is regular, then its eigenvalues are precisely:
  1. a)
    the random eigenvalues of (7) with the same $U$ and $V$;
  2. b)
    the true eigenvalues of $P(\lambda)$.
Proof 3
We know that the true and random eigenvalues of $\widetilde{P}(\lambda)$ are independent of $\tau$ and $B_{0},\dots,B_{k}$. Let $\lambda_{0}$ be such an eigenvalue. Then it follows from Theorem 3.2 that a right eigenvector has the form $V_{\perp}s$ for a nonzero $s\in\mathbb{C}^{n-k}$ or a left eigenvector is of the form $U_{\perp}t$ for a nonzero $t\in\mathbb{C}^{n-k}$ (both statements are true for a true eigenvalue and exactly one of the statements is true for a random eigenvalue). It follows that $P_{22}(\lambda_{0})\,s=0$ or $t^{*}P_{22}(\lambda_{0})=0$. Consequently, $\lambda_{0}$ is an eigenvalue of $P_{22}(\lambda)$. Since (7) altogether has $d(n-k)$ eigenvalues from the groups (a), (b), and (c) from Theorem 3.2, it follows by a simple counting argument that these are all eigenvalues of $P_{22}(\lambda)$$\qed$
Remark 3.6
If we use the slightly weaker concept of genericity with respect to the real and imaginary parts of the entries of $U$ and $V$ in Proposition 3.5, then the regularity of the pencil $U_{\perp}^{*}P(\lambda)\,V_{\perp}$ is generically guaranteed by the results from [11]: since the normal rank of $P(\lambda)$ is $n-k$, it follows that the $(n-k)\times(n-k)$ matrix polynomial $\widetilde{U}^{*}P(\lambda)\widetilde{V}$ is regular, generically with respect to the real and imaginary parts of the entries of $n\times(n-k)$ matrices $\widetilde{U}$ and $\widetilde{V}$. Then using [11, Prop. 9.3] it follows that $U_{\perp}^{*}P(\lambda)\,V_{\perp}$ is regular also generically with respect to the real and imaginary parts of the entries of $U,V\in\mathbb{C}^{n\times k}$.
Another option that can be used to compute the true eigenvalues, is to augment a singular matrix polynomial in such way that it becomes regular while the true eigenvalues remain unchanged. A generalization of the augmented method from [11] is to use the $(n+k)\times(n+k)$ augmented matrix polynomial
$P_{a}(\lambda):=\left[\begin{array}[]{cc}P(\lambda)&UQ_{1}(\lambda)\\ Q_{2}(\lambda)V^{*}&0\end{array}\right],$ (11)
where $Q_{i}(\lambda)=B_{0}^{(i)}+\lambda B_{1}^{(i)}+\cdots+\lambda^{d}B_{d}^{(i)}$ for $i=1,2$ is a regular $k\times k$ matrix polynomial of degree $d$, and $U,V\in\mathbb{C}^{n\times k}$. The following result is a generalization of [11, Prop. 5.1].
Proposition 3.7
Let $P(\lambda)$ be a complex $n\times n$ matrix polynomial of normal rank $n-k$ such that all its eigenvalues are semisimple. Assume that the regular $k\times k$ matrix polynomials $Q_{1}(\lambda)$ and $Q_{2}(\lambda)$ of degree $d$ are chosen in such a way that their $2dk$ eigenvalues are pairwise distinct. Furthermore, let $U,V\in\mathbb{C}^{n\times k}$ have orthonormal columns such that the augmented matrix polynomial (11) is regular and $(U,V^{*})\in\Omega$ holds, where $\Omega$ is the generic set from Theorem 3.2. Then the matrix polynomial (11) has the following eigenvalues:
  1. a)
    $2dk$ prescribed eigenvalues, which are precisely the eigenvalues of $Q_{1}(\lambda)$ and $Q_{2}(\lambda)$;
  2. b)
    the random eigenvalues of (7) with the same $U$ and $V$;
  3. c)
    the true eigenvalues of $P(\lambda)$.
Proof 4
For part a), if $\mu$ is an eigenvalue of $Q_{1}(\lambda)$ with an eigenvector $w\in\mathbb{C}^{k}$, then
$P_{a}(\mu)\left[\begin{array}[]{cc}0\\ w\end{array}\right]=0$
and $\mu$ is an eigenvalue of (11). In a similar way, if $\mu$ is an eigenvalue of $Q_{2}(\lambda)$ with a left eigenvector $z\in\mathbb{C}^{k}$, then
[…]
and $\mu$ is an eigenvalue of (11). As the eigenvalues of $Q_{1}(\lambda)$ and $Q_{2}(\lambda)$ are pairwise distinct, this gives $2dk$ eigenvalues of $P_{a}(\lambda)$.
For b) and c) we consider the perturbed pencil (7) with the same $U$ and $V$ and an arbitrary $Q(\lambda)$ satisfying the hypothesis of Theorem 3.2 exploiting that the random eigenvalues of (7) are independent of $\tau$ and $Q(\lambda)$. Let $U_{\perp}$ and $V_{\perp}$ be matrices whose columns form bases for the orthogonal complement of the ranges of $U$ and $V$, respectively. If $\mu$ is a random eigenvalue of (7) then it follows from Theorem 3.2 that either the corresponding right eigenvector has the form $V_{\perp}s$ for a nonzero $s\in\mathbb{C}^{n-k}$ or the corresponding left eigenvector has the form $U_{\perp}t$ for a nonzero $t\in\mathbb{C}^{n-k}$. Then either
[…]
and $\mu$ is thus an eigenvalue of $P_{a}(\lambda)$.
For c), similarly as in b), we know from Theorem 3.2 that $\mu$ is a true eigenvalue of $P(\lambda)$ when $\mu$ is an eigenvalue of (7) with the right and left eigenvector of the form $V_{\perp}s$ and $U_{\perp}t$, where $s,t\in\mathbb{C}^{n-k}$. It follows that
[…]
therefore $\mu$ is an eigenvalue of $P_{a}(\lambda)$.
As all eigenvalues in c) are semisimple and all eigenvalues in a) and b) are simple, it follows by a counting argument that these are all the eigenvalues of $P_{a}(\lambda)$$\qed$
From the above proof it is easy to see how to extract the true eigenvalues of $P(\lambda)$ from the eigenvalues of $P_{a}(\lambda)$. If $(\theta,x,y)$ is an eigentriplet of $P_{a}(\lambda)$, where
$x=\left[\begin{array}[]{cc}x_{1}\\ x_{2}\end{array}\right],\quad y=\left[\begin{array}[]{cc}y_{1}\\ y_{2}\end{array}\right]$
are in block form in accordance with (11), then $\theta$ is a true eigenvalue if and only if $x_{2}=y_{2}=0$. Here we assume that the matrix polynomials $Q_{1}(\lambda)$ and $Q_{2}(\lambda)$ are chosen in such way that all prescribed eigenvalues differ from the true eigenvalues of $P(\lambda)$.
4 Three methods
In this section we present three numerical methods for singular PEPs that are generalizations of methods for generalized eigenvalue problems from [10] and [11]. All methods return the regular eigenvalues of a singular PEP together with the reciprocals of the condition numbers. Based on the reciprocals of the condition numbers and the gaps between the eigenvalues we can then classify the regular eigenvalues into finite and infinite ones. This is done in Algorithm 4 that is based on the heuristic approach proposed in [11, Alg. 3], for more details see [11, Sec. 6].
The first method is based on Theorem 3.2 and is a generalization of [10, Alg. 1]. In this method we perturb a singular PEP into a regular PEP of the same size. In all algorithms, $\varepsilon$ denotes the machine precision.
Algorithm 1: Eigenvalues of a singular PEP by a rank-completing perturbation.
Input: $A_{0}$, …, $A_{d}\in\mathbb{C}^{n\times n}$, $k=n-\text{nrank}(P)$, perturbation constant $\tau$ (default $10^{-2}$), threshold $\delta$ (default $\varepsilon^{1/2}$).
Output: Eigenvalues of $P(\lambda)=A_{0}+\cdots+\lambda^{d}A_{d}$ with reciprocals of the condition numbers.
1: Select random $n\times k$ matrices $U$ and $V$ with orthonormal columns. 2: Select random $k\times k$ matrices $B_{0},\dots,B_{d}$. 3: Compute the eigenvalues $\lambda_{i}$, $i=1,\dots,dn$, and normalized right and left eigenvectors $x_{i}$ and $y_{i}$ of $\widetilde{P}(\lambda)=P(\lambda)+\tau\,U(B_{0}+\lambda B_{1}+\cdots+\lambda^{d}B_{d})\,V^{*}$. 4: Compute $\alpha_{i}=\|V^{*}x_{i}\|$, $\beta_{i}=\|U^{*}y_{i}\|$, $i=1,\dots,dn$. 5: Compute $\gamma_{i}=|y_{i}^{*}P^{\prime}(\lambda_{i})\,x_{i}|\cdot(1+|\lambda_{i}|^{2}+\cdots+|\lambda_{i}|^{2d})^{-1/2}$ for $i=1,\dots,dn$. 6: Return $\lambda_{i}$ and $\gamma_{i}$ for those $i=1,\dots,dn$ such that $\max(\alpha_{i},\beta_{i})<\delta$.
The second method, which is based on Proposition 3.5, is a generalization of [11, Alg. 1]. In this method we project an $n\times n$ singular PEP of normal rank $n-k$ to an $(n-k)\times(n-k)$ regular PEP.
Algorithm 2: Eigenvalues of a singular PEP by projection.
Input: $A_{0},\dots,A_{d}\in\mathbb{C}^{n\times n}$, $k=n-\text{nrank}(P)$, threshold $\delta$ (default $\varepsilon^{1/2}$).
Output: Eigenvalues of $P(\lambda)=A_{0}+\cdots+\lambda^{d}A_{d}$ with reciprocals of the condition numbers.
1: Select random unitary $n\times n$ matrices $\widehat{W}$ and $\widehat{Z}$, and partition them as $\widehat{W}=[W\ \,W_{\perp}]$ and $\widehat{Z}=[Z\ \,Z_{\perp}]$, where $W$ and $Z$ have $n-k$ columns. 2: Compute the eigenvalues $\lambda_{i}$, $i=1,\dots,d(n-k)$, and normalized right and left eigenvectors $x_{i}$ and $y_{i}$ of $\widetilde{P}(\lambda)=W^{*}A_{0}Z+\lambda\,W^{*}A_{1}Z+\cdots+\lambda^{d}\,W^{*}A_{d}Z$. 3: Compute $\alpha_{i}=\|W_{\perp}^{*}P(\lambda_{i})Zx_{i}\|$, $\beta_{i}=\|y_{i}^{*}W^{*}P(\lambda_{i})\,Z_{\perp}\|$, $i=1,\dots,d(n-k)$. 4: Compute $\gamma_{i}=|y_{i}^{*}W^{*}P^{\prime}(\lambda_{i})\,Zx_{i}|\,(1+|\lambda_{i}|^{2}+\cdots+|\lambda_{i}|^{2d})^{-1/2}$ for $i=1,\dots,d(n-k)$. 5: Return $\lambda_{i}$ and $\gamma_{i}$ for those $i=1,\dots,d(n-k)$ such that $\max(\alpha_{i},\beta_{i})<\delta\,(\,\|A_{0}\|+|\lambda_{i}|\,\|A_{1}\|+\cdots+|\lambda_{i}|^{d}\,\|A_{d}\|\,)$.
The third and final approach, based on Proposition 3.7, is an extension of [11, Alg. 2]. Here we augment an $n\times n$ singular PEP of normal rank $n-k$ to an $(n+k)\times(n+k)$ regular PEP.
Algorithm 3: Eigenvalues of a singular PEP by augmentation.
Input: $A_{0},\dots,A_{d}\in\mathbb{C}^{n\times n}$, $k=n-\text{nrank}$, threshold $\delta$ (default $\varepsilon^{1/2}$).
Output: Eigenvalues of $P(\lambda)=A_{0}+\cdots+\lambda^{d}A_{d}$ with reciprocals of the condition numbers.
1: Select random $n\times k$ matrices $U$ and $V$ with orthonormal columns. 2: Select random $k\times k$ matrices $B_{0}^{(i)},\dots,B_{d}^{(i)}$ for $i=1,2$. 3: Compute the eigenvalues $\lambda_{i}$, $i=1,\dots,d(n+k)$, and normalized right and left eigenvectors $\left[\!\!\begin{array}[]{c}x_{i1}\\ x_{i2}\end{array}\!\!\right]$ and $\left[\!\!\begin{array}[]{c}y_{i1}\\ y_{i2}\end{array}\!\!\right]$ of the augmented matrix polynomial (11). 4: Compute $\alpha_{i}=\|x_{i2}\|$,  $\beta_{i}=\|y_{i2}\|$,  $i=1,\dots,d(n+k)$. 5: Compute $\gamma_{i}=|y_{i1}^{*}P^{\prime}(\lambda_{i})\,x_{i1}|\cdot(1+|\lambda_{i}|^{2}+\cdots+|\lambda_{i}|^{2d})^{-1/2}$,  $i=1,\ldots,d(n+k)$. 6: Return $\lambda_{i}$ and $\gamma_{i}$ for those $i=1,\ldots,d(n+k)$ such that $\max(\alpha_{i},\beta_{i})<\delta$.
To classify a computed regular eigenvalue $\lambda_{i}$ as finite or infinite, we use the computed reciprocal of the condition number $\gamma_{i}$ combined with the relative gap of an eigenvalue, which is defined as
${\rm gap}_{i}=\min_{j\neq i}\frac{|\lambda_{j}-\lambda_{i}|}{(1+|\lambda_{i}|^{2})^{1/2}}.$
The criterion is based on the fact that $y_{i}^{*}P^{\prime}(\lambda_{i})\,x_{i}\neq 0$ if $\lambda_{i}$ is a simple finite eigenvalue. A singular PEP can also have multiple eigenvalues for which $\gamma_{i}$ might be 00, yet in practice, as explained in more details in [11], an eigenvalue of multiplicity $m$ is numerically evaluated as $m$ simple eigenvalues with nonzero values $\gamma_{i}$ and this way multiple finite eigenvalues are properly identified in most cases. We use the relative gap to further improve the detection when $\gamma_{i}$ is small. We expect that a computed representative $\lambda_{i}$ of a multiple finite eigenvalue will have a small $\gamma_{i}$ but also a small ${\rm gap}_{i}$. On the other hand, a (multiple) infinite eigenvalue will usually appear as a finite eigenvalue with $\gamma_{i}$ very close to zero and a large ${\rm gap}_{i}$ (close to $1$).
Algorithm 4 is based on [11, Alg. 3], with slightly different default values. Its input are regular eigenvalues and reciprocals of condition numbers that we get from any of Algorithms 1–3.
Algorithm 4: Extraction of finite eigenvalues from the regular eigenvalues of a singular PEP.
Input: Eigenvalues $\lambda_{i}$ and reciprocals $\gamma_{i}$ of condition numbers, $i=1,\ldots,r$, for regular eigenvalues of PEP, thresholds $\delta_{1}$ (default $\varepsilon$), $\delta_{2}$ (default $10^{4}\,\varepsilon$), and $\xi_{2}$ (default 0.01).
Output: Finite eigenvalues.
1: Compute ${\rm gap}_{i}=\min_{j\neq i}|\lambda_{j}-\lambda_{i}|\cdot(1+|\lambda_{i}|^{2})^{-1/2}$,  $i=1,\ldots,r$. 2: If $\gamma_{i}<\delta_{1}$,  $i=1,\ldots,r$, flag $\lambda_{i}$ as an infinite eigenvalue. 3: If $\gamma_{i}<\delta_{2}$ and ${\rm gap}_{i}>\xi_{2}$,  $i=1,\ldots,r$, flag $\lambda_{i}$ as an infinite eigenvalue. 4: Return $\lambda_{i}$ for those $i=1,\ldots,r$ such that $\lambda_{i}$ is not flagged as an infinite eigenvalue.
5 Comparison to linearization
A common way to solve a PEP is to apply one of the numerous possible linearizations. If a linearization is strong, then from a singular matrix polynomial $P(\lambda)$ we get a singular matrix pencil $L(\lambda):=E+\lambda F$ such that the finite and infinite eigenvalues of $L$ and $P$ coincide. Suppose that $P$ is an $n\times n$ matrix polynomial of degree $d$ and normal rank $n-k$. We can apply the following two procedures to compute the eigenvalues of $P$.
  1. 1)
    First, apply Algorithm 2 and project $P(\lambda)$ into a regular $(n-k)\times(n-k)$ matrix polynomial $\widetilde{P}(\lambda)$ of degree $d$. Then, to compute the eigenvalues and eigenvectors of $\widetilde{P}$, linearize it into a regular matrix pencil $L_{1}(\lambda)$ of size $d(n-k)\times d(n-k)$.
  2. 2)
    First, linearize $P(\lambda)$ into a singular matrix pencil $L(\lambda)$ of size $dn\times dn$. Then, apply Algorithm 2 to $L$ and project it into a regular matrix pencil $L_{2}(\lambda)$.
Suppose for example that in 2) we use a strong linearization from the space $\mathbb{L}_{1}(P)$ introduced in [17]. If $\varepsilon_{1}\leq\cdots\leq\varepsilon_{k}$ are the right minimal indices of $P(\lambda)$, then by [3, Thm. 5.10] the right minimal indices of $L(\lambda)$ are $\varepsilon_{1}+d-1\leq\cdots\leq\varepsilon_{k}+d-1$, while the left minimal indices of $P(\lambda)$ and $L(\lambda)$ are equal. It then follows from the index sum theorem that the normal rank of $L(\lambda)$ is $nd-k$ and thus $L_{2}(\lambda)$ is of size $(nd-k)\times(nd-k)$. The results in [3] show that also types of strong linearizations other than from the space $\mathbb{L}_{1}(P)$ will lead to projected pencils of the same size.
This shows that the two approaches are not equivalent and if we apply 1) we end up with a smaller regular pencil. This requires less computational work, in particular when $d$ is large and the normal rank of $P(\lambda)$ is small. Also, in both combinations we have to determine the normal rank at a certain point, which we compute from the rank of $P(\lambda)$ or $L(\lambda)$ evaluated at random values $\lambda$. While in 1) we compute it from a rank of an $n\times n$ matrix, in 2) we have to work with $nd\times nd$ matrices.
Based on the above, it seems to be more efficient to first project the polynomial into a regular one and then linearize it.
6 Applications and numerical examples
We test our numerical methods for singular PEPs on problems coming from several applications. Two motivating problems are certain bivariate polynomial systems and ZGV points, which may not yet be widely known. We also treat some problems from [15] and other papers. In addition, although there are additional references that mention the need of solving singular PEPs, they usually do not provide concrete examples. We hope that this paper may also encourage new research that will lead to additional applications that involve eigenvalues of singular matrix polynomials.
All numerical experiments have been carried out in MATLAB R2023a using standard double precision. To compute eigenvalues and eigenvectors of PEPs in Algorithms 1–3 we use polyeig in Matlab. We modified the method so that it returns the left eigenvectors (which are computed in the original polyeig for the estimation of condition numbers but are not returned) as well. For QEPs we use quadeig from [9], which is more efficient and accurate than polyeig and returns both right and left eigenvectors.
Inspired by [15] we performed $N=10000$ runs of Algorithms 1–3, followed by Algorithm 4, for each of the problems and counted the number of successful runs $M$ when the method returned the expected number of finite eigenvalues. The ratio $p=M/N$ is the empirical probability that the algorithm returns a correct answer. For problems, for which we know the exact solution, we also computed the maximal absolute error of the computed eigenvalues (in all successful attempts) and we report this as the maximal error. We used the threshold parameter $\delta=10^{-10}$ for Algorithm 1 and $\delta=10^{-12}$ for Algorithms 2 and 3. In Algorithm 4 we used the default values for $\delta_{1},\delta_{2}$, and $\xi_{2}$. Instead of fine-tuning the parameters for each of the examples, we have selected the optimal overall settings for all the examples.
6.1 Polynomial equations
Consider a system of equations involving bivariate polynomials of the form
$p(\lambda^{2},\mu)=0,\qquad q(\lambda,\mu)=0,$
where in all monomials of the first polynomial the degree of $\lambda$ is even. The uniform determinantal representations of [1] render $A_{i}$, $B_{i}$, $C_{i}$, for $i=1,2$, such that
$p(\lambda^{2},\mu)$ $=\det(A_{1}+\lambda^{2}\,B_{1}+\mu\,C_{1}),$
$q(\lambda,\mu)$ $=\det(A_{2}+\lambda\,B_{2}+\mu\,C_{2}),$
and the associated two-parameter eigenvalue problem
$(A_{1}+\lambda^{2}\,B_{1}+\mu\,C_{1})\,x$ $=0,$
$(A_{2}+\lambda\,B_{2}+\mu\,C_{2})\,y$ $=0.$
If we take the Kronecker product of the first equation and $C_{2}y$ and subtract the Kronecker product of $C_{1}x$ and the second equation, we obtain the QEP
$(A_{1}\otimes C_{2}-C_{1}\otimes A_{2})\,z-\lambda\,(C_{1}\otimes B_{2})\,z+\lambda^{2}\,(B_{1}\otimes C_{2})\,z=0,$ (12)
where $z=x\otimes y$. This QEP is typically singular. We give a simple example.
Example 6.1
Consider the polynomial system
$p(\lambda^{2},\mu)$ $=1+2\lambda^{2}+3\mu+4\lambda^{4}+5\lambda^{2}\mu+6\mu^{2},$
$q(\lambda,\mu)$ $=6+5\lambda+4\mu+3\lambda^{2}+2\lambda\mu+\mu^{2},$
which has 8 finite solutions
$(\lambda_{1,2},\,\mu_{1,2})$ $=(-0.658067\pm 0.750641i,\ -0.816143\pm 0.414507i),$
$(\lambda_{3,4},\,\mu_{3,4})$ $=(-1.332648\pm 0.355433i,\ -0.287759\pm 1.672047i),$
$(\lambda_{5,6},\,\mu_{5,6})$ $=(\phantom{-}0.475211\pm 1.902116i,\ -0.097155\mp 3.062254i),$
$(\lambda_{7,8},\,\mu_{7,8})$ $=(\phantom{-}2.765503\pm 0.580944i,\ -5.548943\pm 3.891223i).$
Then
[…]
The corresponding QEP (12) is
$P(\lambda)={\mbox{\scriptsize$\left[\!\!\begin{array}[]{ccccccccc}0&0&0&0&0&0&0&0&0\\ 0&0&0&2+\lambda^{2}&0&0&1&0&0\\ 0&0&0&0&-2-4\lambda^{2}&0&0&-1&0\\ 0&-30-18\lambda&-6&0&0&0&0&0&0\\ -21-12\lambda+5\lambda^{2}&-36&6\lambda&1&0&0&-\lambda^{2}&0&0\\ -6&-3-5\lambda^{2}&0&0&-1&0&0&\lambda^{2}&0\\ 0&0&0&0&5+3\lambda&1&0&0&0\\ 1&0&0&4+2\lambda&6&-\lambda&0&0&0\\ 0&-1&0&1&0&0&0&0&0\end{array}\!\!\right]$}},$ (13)
with normal rank equal to $8$ so that $P(\lambda)$ is a singular QEP. If we apply Algorithm 2, we obtain the finite eigenvalues in Table 1 which agree with the $\lambda$-components of the solution.
We run Algorithms 1–3 for 10000 times on (13) and computed the empirical probability of success $p_{i}$ and maximal error ${\rm maxerr}_{i}$ for $i=1,2,3$. The results are $p_{1}=p_{2}=p_{3}=1$, ${\rm maxerr}_{1}=1.0\cdot 10^{-11}$, ${\rm maxerr}_{2}=7.6\cdot 10^{-11}$, and ${\rm maxerr}_{3}=1.3\cdot 10^{-10}$.
Table 1: Results of Algorithm 2 (projection to normal rank) applied to the singular QEP (13) with matrices from Example 6.1.
$j$ $ik_{j}$ $\quad\ \gamma_{j}$ $\quad\ \alpha_{j}$ $\quad\ \beta_{j}$ ${\rm gap}_{j}$ Type
1 $-0.658067+0.750641i$ $3.5\cdot 10^{-2}$ $7.5\cdot 10^{-17}$ $4.2\cdot 10^{-17}$ $0.24$ Finite true
2 $-0.658067-0.750641i$ $1.8\cdot 10^{-2}$ $3.9\cdot 10^{-17}$ $3.9\cdot 10^{-17}$ $0.24$ Finite true
3 $-1.332648+0.355434i$ $9.1\cdot 10^{-3}$ $6.5\cdot 10^{-17}$ $3.7\cdot 10^{-17}$ $0.21$ Finite true
4 $-1.332648-0.355434i$ $4.6\cdot 10^{-3}$ $2.7\cdot 10^{-17}$ $2.5\cdot 10^{-17}$ $0.21$ Finite true
5 $\phantom{-}0.475211+1.902116i$ $3.2\cdot 10^{-3}$ $4.0\cdot 10^{-17}$ $2.3\cdot 10^{-17}$ $0.44$ Finite true
6 $\phantom{-}0.475211-1.902116i$ $2.6\cdot 10^{-3}$ $5.4\cdot 10^{-17}$ $2.8\cdot 10^{-17}$ $0.44$ Finite true
7 $\phantom{-}2.765503+0.580944i$ $9.0\cdot 10^{-4}$ $1.5\cdot 10^{-17}$ $1.6\cdot 10^{-17}$ $0.28$ Finite true
8 $\phantom{-}2.765503-0.580944i$ $8.8\cdot 10^{-4}$ $6.1\cdot 10^{-17}$ $1.6\cdot 10^{-17}$ $0.28$ Finite true
9 $\phantom{-}2.19328\cdot 10^{7}+2.41952\cdot 10^{6}i$ $8.8\cdot 10^{-24}$ $9.5\cdot 10^{-17}$ $9.7\cdot 10^{-17}$ $0.35$ Infinite true
10 $-2.19328\cdot 10^{7}-2.41952\cdot 10^{6}i$ $4.6\cdot 10^{-24}$ $1.1\cdot 10^{-16}$ $1.2\cdot 10^{-16}$ $0.35$ Infinite true
11 $\phantom{-}1.91184\cdot 10^{7}-4.69967\cdot 10^{6}i$ $7.3\cdot 10^{-24}$ $1.7\cdot 10^{-16}$ $2.6\cdot 10^{-17}$ $0.39$ Infinite true
12 $-1.91184\cdot 10^{7}+4.69967\cdot 10^{6}i$ $4.3\cdot 10^{-24}$ $1.8\cdot 10^{-16}$ $7.8\cdot 10^{-17}$ $0.39$ Infinite true
13 $\infty$ $0.0$ $9.6\cdot 10^{-17}$ $1.3\cdot 10^{-16}$ $1.00$ Infinite true
14 $\infty$ $0.0$ $1.5\cdot 10^{-16}$ $5.6\cdot 10^{-17}$ $1.00$ Infinite true
15 $\infty$ $0.0$ $5.6\cdot 10^{-17}$ $6.2\cdot 10^{-17}$ $1.00$ Infinite true
16 $\infty$ $0.0$ $1.4\cdot 10^{-17}$ $5.9\cdot 10^{-17}$ $1.00$ Infinite true
In the following table we give the generic normal ranks corresponding to some low values of the total degree $d$ (in $\lambda^{2}$ and $\mu$ for $p$; in $\lambda$ and $\mu$ for $q$).
Degree $d$ $n=(2d-1)^{2}$ nrank Missing
2 119 18 11
3 125 21 14
4 149 40 19
5 181 65 16
6 121 96 25
For $n=(2d-1)^{2}$, the normal rank seems to be $\text{nrank}=d\,(3d-2)$, with missing rank $(d-1)^{2}$. Therefore, this is a problem of dimension $\sim 4d^{2}$, with normal rank $\sim 3d^{2}$ and rank-completing dimension $\sim d^{2}$.
6.2 ZGV points
In the study of anisotropic elastic waveguides (see, e.g., [19]) we obtain an eigenvalue problem
$\big{(}(ik)^{2}L_{2}+ikL_{1}+L_{0}+\omega^{2}M\big{)}\,u=0,$ (14)
where $L_{0}$, $L_{1}$, $L_{2}$, $M$ are real $n\times n$ matrices obtained by a discretization of a boundary value problem. The solution are dispersion curves $\omega=\omega(k)$, where the angular frequency $\omega$ of a guided wave is related to the wavenumber $k$. We are interested in the zero-group-velocity (ZGV) points on dispersion curves, where $\omega$ and $k$ are real and $\omega^{\prime}(k)=0$. If we assume that $u=u(k)$ and $\omega(k)$ are differentiable, we obtain from (14) (see [14] for details) that at a ZGV point $(\omega,k)$ we have
$\big{(}(ik)^{2}\,\widetilde{L}_{2}+ik\,\widetilde{L}_{1}+\widetilde{L}_{0}+\omega^{2}\,\widetilde{M}\big{)}\,\widetilde{u}=0,$ (15)
where
$\widetilde{L}_{2}=\left[\begin{matrix}L_{2}&0\\ 0&L_{2}\end{matrix}\right],\quad\widetilde{L}_{1}=\left[\begin{matrix}L_{1}&0\\ 2L_{2}&L_{1}\end{matrix}\right],\quad\widetilde{L}_{0}=\left[\begin{matrix}L_{0}&0\\ L_{1}&L_{0}\end{matrix}\right],\quad\widetilde{M}=\left[\begin{matrix}M&0\\ 0&M\end{matrix}\right],\quad\widetilde{u}=\left[\begin{matrix}u\\ u^{\prime}\end{matrix}\right].$
Equations (14) and (15) form a quadratic two-parameter eigenvalue problem [18, 12]. If we take the Kronecker product of (14) and $\widetilde{M}\widetilde{u}$, and subtract the Kronecker product of $Mu$ and (15), then we get a $2n^{2}\times 2n^{2}$ QEP
$\big{(}(ik)^{2}\,\Gamma_{2}+ik\,\Gamma_{1}+\Gamma_{0}\big{)}\,z=0,$ (16)
where $z=u\otimes\widetilde{u}$ and
$\Gamma_{2}=L_{2}\otimes\widetilde{M}-\widetilde{M}\otimes L_{2},\quad\Gamma_{1}=L_{1}\otimes\widetilde{M}-\widetilde{M}\otimes L_{1},\quad\Gamma_{0}=L_{0}\otimes\widetilde{M}-\widetilde{M}\otimes L_{0}.$
With respect to the entries of the matrices $L_{0}$, $L_{1}$, $L_{2}$, $M$, the normal rank of the quadratic matrix polynomial (16) generically is $2n^{2}-n$ and hence the problem is singular.
Example 6.2
For a simple example we take
$L_{2}={\mbox{\scriptsize$\left[\!\!\begin{array}[]{rr}1&1\\ 1&2\end{array}\!\!\right]$}},\quad L_{1}={\mbox{\scriptsize$\left[\!\!\begin{array}[]{rr}0&3\\ -3&0\end{array}\!\!\right]$}},\quad L_{0}={\mbox{\scriptsize$\left[\!\!\begin{array}[]{rr}-2&1\\ 1&-2\end{array}\!\!\right]$}},\quad M={\mbox{\scriptsize$\left[\!\!\begin{array}[]{rr}3&1\\ 1&4\end{array}\!\!\right]$}}.$
We selected the matrices so that $L_{0}$ is symmetric, $L_{1}$ is skew-symmetric, and $L_{2},M$ are symmetric positive definite. This way the matrices have the same properties as the larger matrices in [14], where ZGV points of antisymmetric Lamb waves in an austenitic steel plate are computed.
The corresponding quadratic matrix polynomial (16) with $8\times 8$ matrices $\Gamma_{2},\Gamma_{1},\Gamma_{0}$ has normal rank $6$. The problem has a double eigenvalue at 00 of geometric multiplicity 2 because $\mathop{\rm rank}\nolimits(\Gamma_{0})=4$. Algorithm 1 returns the eigenvalues displayed in Table 2. Notice that the double eigenvalue 00 is computed exactly due to a preprocessing step in quadeig that reveals and deflates the zero and infinite eigenvalues contributed by singular leading and trailing matrix coefficients. Notice also that, although 00 is a double eigenvalue, the corresponding reciprocals of the condition numbers $\gamma_{1}$ and $\gamma_{2}$ are far away from zero because the eigenvalue is semisimple.
There are four eigenvalues $ik$ on the imaginary axis that give $k$-coordinates of ZGV points $(\omega,k)$. For each of these values of $k$ we solve the generalized eigenvalue problem (14) to compute $\omega$ and a corresponding eigenvector $u$. Since all matrices are symmetric, $u$ is also a left eigenvector and the condition for a ZGV point is
$u^{*}\,(2ikL_{2}+L_{1})\,u=0,$
see [14]. We get four real ZGV points $(1.110602,0)$, $(0.470226,0)$, $(0.364791,\pm 1.016018)$ which are shown together with real dispersion curves in Figure 1.
As in the previous example, based on 10000 runs, we get $p_{1}=p_{2}=p_{3}=1$, ${\rm maxerr}_{1}=1.7\cdot 10^{-12}$, ${\rm maxerr}_{2}=1.2\cdot 10^{-11}$, and ${\rm maxerr}_{2}=7.0\cdot 10^{-12}$.
Table 2: Results of Algorithm 1 (rank-completing perturbation) applied to the singular QEP (16) with matrices from Example 6.2
$j$ $ik_{j}$ $\quad\ \gamma_{j}$ $\quad\ \alpha_{j}$ $\quad\ \beta_{j}$ ${\rm gap}_{j}$ Type
1 00 $6.7\cdot 10^{-2}$ $1.6\cdot 10^{-15}$ $2.6\cdot 10^{-15}$ $0.00$ Finite true
2 00 $6.3\cdot 10^{-2}$ $1.8\cdot 10^{-15}$ $1.6\cdot 10^{-15}$ $0.00$ Finite true
3 $-4.007967\cdot 10^{-16}+1.016018i$ $6.1\cdot 10^{-2}$ $1.6\cdot 10^{-14}$ $1.3\cdot 10^{-14}$ $0.37$ Finite true
4 $\phantom{-}1.716439\cdot 10^{-15}-1.016018i$ $4.7\cdot 10^{-2}$ $6.5\cdot 10^{-15}$ $4.2\cdot 10^{-14}$ $0.33$ Finite true
5 $-4.004034-7.895649\cdot 10^{-15}i$ $5.3\cdot 10^{-3}$ $1.2\cdot 10^{-14}$ $5.3\cdot 10^{-14}$ $0.61$ Finite true
6 $\phantom{-}4.004034-1.035145\cdot 10^{-14}i$ $4.4\cdot 10^{-3}$ $1.8\cdot 10^{-14}$ $2.5\cdot 10^{-14}$ $0.87$ Finite true
7 $\infty$- ∞ $0.0$ $4.2\cdot 10^{-15}$ $4.8\cdot 10^{-15}$ $1.00$ Infinite true
[…]
9 $-0.202213-0.301941i$ $3.7\cdot 10^{-3}$ $5.3\cdot 10^{-15}$ $0.54$ $0.18$ Random right
10 $-3.322434+8.098093i$ $1.2\cdot 10^{-4}$ $8.2\cdot 10^{-15}$ $0.56$ $0.72$ Random right
11 $\phantom{-}0.114927+0.193407i$ $1.1\cdot 10^{-2}$ $0.29$ $8.6\cdot 10^{-15}$ $0.11$ Random left
12 $\phantom{-}9.990344-1.667635\cdot 10^{1}i$ $4.7\cdot 10^{-5}$ $0.23$ $2.4\cdot 10^{-14}$ $0.91$ Random left
13 $-0.815756+2.188693i$ $3.2\cdot 10^{-3}$ $0.24$ $0.72$ $0.39$ Prescribed
14 $-0.815756-2.188693i$ $3.1\cdot 10^{-3}$ $0.29$ $0.60$ $0.39$ Prescribed
15 $-1.552148-1.225251i$ $9.4\cdot 10^{-4}$ $0.17$ $0.55$ $0.43$ Prescribed
16 $-1.552148+1.225251i$ $7.9\cdot 10^{-4}$ $0.15$ $0.52$ $0.43$ Prescribed
Figure 1: Real dispersion curves $\omega(k)$ and ZGV points of Example 6.2
6.3 Set of singular QEP examples
In a recent paper [15], Kressner and Šain Glibić construct several examples of singular QEPs to test their numerical method, where they perturb a singular QEP into a regular problem by a small random full rank perturbation, compute the eigenvalues together with the left and right eigenvectors and then identify the finite eigenvalues based on the size of the condition number (6). This might be compared to Algorithm 1 followed by Algorithm 4, but since a full rank perturbations are used in [15], the eigenvalues of the perturbed problem are perturbed as well. We denote the algorithm from [15] as Algorithm KŠG.
For each of the problems they perform 1000 runs of the algorithm with different random perturbation matrices and computed empirical probabilities as the ratio of runs when the algorithm found all the finite eigenvalues. We tested Algorithm KŠG and Algorithms 1–3, followed by Algorithm 4, on the most challenging examples from [15] in a similar way. As another option we tested standard linearizations from $\mathbb{L}_{1}(P)$ combined with the projection algorithm for a singular generalized eigenvalue problem from [11], this option is denoted Linearization.
Example 6.3
Our selected QEPs are (the first four problems are [15, Exs. 5–8], while 5) is new):
  1. 1)
    $Q(\lambda)=Z^{*}(K+\lambda C+\lambda^{2}M)\,W$ with $8\times 8$ matrices and finite eigenvalues $\lambda_{j}=1+10^{-5}j$ for $j=1,\dots,5$. The nonzero elements of $K,C,M$ are $m_{j,j+1}=c_{jj}=1$ and $c_{j,j+1}=k_{jj}=-\lambda_{j}$ for $j=1,\dots,5$, W = orth(rand(8)) and Z = orth(rand(8)).
  2. 2)
    $Q(\lambda)=Z^{*}(K+\lambda C+\lambda^{2}M)\,W$ with $11\times 1$ 11 matrices and finite eigenvalues $\lambda_{1}=0$, $\lambda_{j}=1/j$ for $j=2,\dots,8$. The nonzero elements of $K,C,M$ are $m_{j,j+1}=c_{jj}=1$ and $c_{j,j+1}=k_{jj}=-\lambda_{j}$ for $j=1,\dots,8$, W = orth(rand(11)) and Z = orth(rand(11)).
  3. 3)
    The reversal of 2), i.e., $Q(\lambda)=Z^{*}(M+\lambda C+\lambda^{2}K)\,W$. Finite eigenvalues are $\lambda_{j}=j+1$ for $j=1,\dots,7$.
  4. 4)
    $Q(\lambda)=Z^{*}(DMD+\lambda DCD+\lambda^{2}DKD)\,W$, where $D$ is the scaling matrix $D=\mathop{\rm diag}\nolimits(1,a^{2},a,1,a^{3},1,a^{4},a^{5},a^{6},1,1)$ for $a=2,4,6,8$. The other matrices and finite eigenvalues are the same as in 3). The diagonal scaling makes the eigenvalues more ill conditioned and more difficult to detect and compute accurately.
  5. 5)
    $Q(\lambda)=Z^{*}(D^{-1}KD+\lambda D^{-1}CD+\lambda^{2}D^{-1}MD)\,W$ with $8\times 8$ matrices $K,C,M$, whose nonzero elements are $m_{j,j+1}=c_{jj}=1$, $m_{13}=m_{24}=1$, and $c_{j,j+1}=k_{jj}=-1$ for $j=1,\dots,5$, and $D=\mathop{\rm diag}\nolimits(1,a^{3},a^{6},a^{2},a^{5},a,a^{4},a^{7})$ for $a=1,2,3$. The problem has finite eigenvalue $1$ with algebraic multiplicity $4$ and geometric multiplicity $3$.
For each problem we carried out $N=10000$ runs of each algorithm and counted the number of runs $F$ when the method failed to return the exact number of eigenvalues. The ratio $p=1-F/N$ is the empirical probability that the algorithm returns the correct answer. We also computed the maximal absolute error of the computed eigenvalues in all successful attempts, we report this as the maximal error. The results are listed in Table 3. For Algorithm KŠG we used parameters $\varepsilon=10^{-12}$ for the size of the perturbation and ${\sf tol}=10^{10}$ for the tolerance. With these settings we obtained better empirical probabilities and smaller error than reported in [15]. For Linearization we used the threshold parameter $\delta=10^{-12}$.
Table 3: Results of Algorithm KŠG [15, Alg. 2], Algorithm 1 (rank-completing perturbation), Algorithm 2 (projection to normal rank), Algorithm 3 (augmentation), and Linearization ([11, Alg. 2] applied to an $\mathbb{L}_{1}$ linearization) applied to the singular QEPs from Example 6.3.
Algorithm KŠG Algorithm 1 Algorithm 2 Algorithm 3 Linearization
Problem $F$ max error $F$ max error $F$ max error $F$ max error $F$ max error
1) 00 $1.5\cdot 10^{-10}$ 00 $6.3\cdot 10^{-13}$ 00 $1.5\cdot 10^{-13}$ 00 $2.5\cdot 10^{-13}$ 00 $8.6\cdot 10^{-13}$
2) 00 $4.8\cdot 10^{-11}$ 00 $1.7\cdot 10^{-13}$ 00 $3.4\cdot 10^{-14}$ 00 $2.0\cdot 10^{-13}$ 00 $4.8\cdot 10^{-13}$
3) 00 $4.8\cdot 10^{-9}$ 00 $4.8\cdot 10^{-12}$ 00 $8.8\cdot 10^{-13}$ 00 $1.6\cdot 10^{-11}$ 00 $6.3\cdot 10^{-12}$
4) $(a=2)$ $1$ $2.5\cdot 10^{-7}$ 00 $9.0\cdot 10^{-11}$ 00 $6.7\cdot 10^{-11}$ 00 $2.2\cdot 10^{-10}$ 00 $5.1\cdot 10^{-10}$
4) $(a=4)$ $3$ $1.6\cdot 10^{-4}$ 00 $6.6\cdot 10^{-8}$ 00 $1.4\cdot 10^{-7}$ 00 $2.0\cdot 10^{-7}$ 00 $6.1\cdot 10^{-7}$
4) $(a=6)$ $68$ $2.6\cdot 10^{-3}$ 00 $9.3\cdot 10^{-6}$ 00 $1.0\cdot 10^{-5}$ 00 $2.1\cdot 10^{-5}$ 00 $1.7\cdot 10^{-5}$
4) $(a=8)$ $1204$ $1.6\cdot 10^{-2}$ $3$ $1.9\cdot 10^{-4}$ $1$ $1.5\cdot 10^{-4}$ $3$ $3.4\cdot 10^{-4}$ $20$ $6.5\cdot 10^{-4}$
5) $(a=1)$ 00 $5.7\cdot 10^{-6}$ 00 $2.8\cdot 10^{-7}$ 00 $3.9\cdot 10^{-7}$ 00 $3.2\cdot 10^{-7}$ 00 $4.4\cdot 10^{-7}$
5) $(a=2)$ $50$ $5.1\cdot 10^{-5}$ 00 $2.9\cdot 10^{-6}$ 00 $2.5\cdot 10^{-6}$ 00 $5.7\cdot 10^{-6}$ 00 $7.2\cdot 10^{-6}$
5) $(a=3)$ $3020$ $1.4\cdot 10^{-4}$ 00 $2.4\cdot 10^{-5}$ 00 $1.4\cdot 10^{-5}$ 00 $3.5\cdot 10^{-5}$ 00 $8.1\cdot 10^{-5}$
The singular QEP in 5) has a quadruple eigenvalue $\lambda=1$. Table 4 shows the output of Algorithm 2 for $a=3$. We can see that the values $\gamma_{i}$ for the quadruple eigenvalue are close to 00 as are the relative gaps, as expected by Algorithm 4.
Table 4: Results of Algorithm 2 (projection to normal rank) applied to the singular QEP 5) from Example 6.3 for $a=3$.
$j$ $\lambda_{j}$ $\quad\ \gamma_{j}$ $\quad\ \alpha_{j}$ $\quad\ \beta_{j}$ ${\rm gap}_{j}$ Type
1 $\phantom{-}1.000000+1.036016\cdot 10^{-15}i$ $1.1\cdot 10^{-4}$ $1.4\cdot 10^{-16}$ $7.7\cdot 10^{-17}$ $2.1\cdot 10^{-11}$ Finite true
2 $\phantom{-}1.000000+1.549034\cdot 10^{-11}i$ $1.3\cdot 10^{-8}$ $1.8\cdot 10^{-16}$ $6.6\cdot 10^{-17}$ $2.1\cdot 10^{-11}$ Finite true
3 $\phantom{-}0.999999-6.066428\cdot 10^{-7}i$ $7.2\cdot 10^{-13}$ $1.4\cdot 10^{-16}$ $4.0\cdot 10^{-17}$ $6.4\cdot 10^{-7}$ Finite true
4 $\phantom{-}1.000000+6.066328\cdot 10^{-7}i$ $7.2\cdot 10^{-13}$ $1.3\cdot 10^{-16}$ $4.3\cdot 10^{-17}$ $6.4\cdot 10^{-7}$ Finite true
5 $-0.748612+7.673687\cdot 10^{-2}i$ $2.0\cdot 10^{-7}$ $1.0\cdot 10^{-16}$ $9.0\cdot 10^{-6}$ $1.18$ Random right
6 $-0.355925-0.716723i$ $1.0\cdot 10^{-7}$ $1.6\cdot 10^{-16}$ $5.2\cdot 10^{-6}$ $1.11$ Random right
7 $\phantom{-}0.591362+0.242587i$ $7.6\cdot 10^{-8}$ $1.3\cdot 10^{-16}$ $2.4\cdot 10^{-6}$ $0.74$ Random right
8 $-6.082501\cdot 10^{-2}+0.828186i$ $6.9\cdot 10^{-8}$ $1.2\cdot 10^{-16}$ $3.9\cdot 10^{-6}$ $1.05$ Random right
9 $\phantom{-}0.579022-0.389633i$ $6.2\cdot 10^{-8}$ $1.3\cdot 10^{-16}$ $2.5\cdot 10^{-6}$ $0.82$ Random right
10 $\phantom{-}3.012867-1.129741\cdot 10^{-2}i$ $2.6\cdot 10^{-8}$ $1.5\cdot 10^{-16}$ $1.3\cdot 10^{-7}$ $0.67$ Random right
6.4 Other applications
Although there are several references to applications of singular matrix polynomials in the literature, we managed to collect just few concrete examples.
Example 6.4
In [13], where a generalization of the Moore–Penrose inverse has been introduced for singular matrix polynomials, the following singular polynomial matrix of degree $5$ is mentioned [13, Ex. 1]:
$P(\lambda)={\mbox{\scriptsize$\left[\!\!\begin{array}[]{ccc}1+4\lambda+5\lambda^{2}+2\lambda^{3}&-1-3\lambda-4\lambda^{2}-3\lambda^{3}-\lambda^{4}&-\lambda-2\lambda^{2}-\lambda^{3}\\[1.42262pt] -1-2\lambda+2\lambda^{2}+5\lambda^{3}+2\lambda^{4}&1+\lambda-\lambda^{2}-3\lambda^{3}-3\lambda^{4}-\lambda^{5}&\lambda-2\lambda^{3}-\lambda^{4}\\[1.42262pt] -1-2\lambda+\lambda^{2}+2\lambda^{3}&1+\lambda-\lambda^{3}-\lambda^{4}&\lambda-\lambda^{3}\end{array}\!\!\right]$}},$ (17)
where $\mathop{\rm nrank}\nolimits(P(\lambda))=1$. The problem has one finite eigenvalue $\lambda=-1$, which is identified correctly by Algorithm 2; results of a typical run are presented in Table 5.
Table 5: Results of Algorithm 2 (projection to normal rank) applied to the singular PEP (17).
$j$ $\lambda_{j}$ $\quad\ \gamma_{j}$ $\quad\ \alpha_{j}$ $\quad\ \beta_{j}$ ${\rm gap}_{j}$ Type
1 $-1.000000+1.016777\cdot 10^{-16}i$ $6.4\cdot 10^{-2}$ $1.3\cdot 10^{-16}$ $1.4\cdot 10^{-16}$ $0.56$ Finite true
2 $-0.734887-0.835476i$ $8.8\cdot 10^{-2}$ $1.2\cdot 10^{-16}$ $5.1\cdot 10^{-2}$ $0.54$ Random right
3 $-1.108317+1.038442i$ $5.5\cdot 10^{-2}$ $1.1\cdot 10^{-16}$ $5.1\cdot 10^{-2}$ $0.48$ Random right
4 $\phantom{-}0.230639-1.285601i$ $2.1\cdot 10^{-1}$ $2.5\cdot 10^{-1}$ $2.7\cdot 10^{-16}$ $0.60$ Random left
5 $-0.190740+0.980385i$ $1.6\cdot 10^{-1}$ $1.7\cdot 10^{-1}$ $9.2\cdot 10^{-17}$ $0.59$ Random left
Note that in this case the projected polynomial is of size $1\times 1$, thus we need to find roots of a scalar polynomial of degree $5$. The corresponding left and right eigenvectors from Algorithm 2 are scalars $x_{i}=y_{i}=1$. Still, together with the initial unitary transformations $U$ and $V$ they provide enough information that we can classify the eigenvalues. If we instead first linearize $P(\lambda)$ into a linear pencil, then we get a singular pencil with matrices $15\times 15$ that has normal rank $13$. This example shows that it may be considerably more efficient to apply Algorithm 2, in particular for polynomials of high degree.
We perform $N=10000$ runs and obtained empirical probabilities $p_{1}=0.9983$, $p_{2}=1$ and $p_{3}=0.9968$ for Algorithms 1–3 with maximal errors ${\rm maxerr}_{1}=7.1\cdot 10^{-13}$, ${\rm maxerr}_{2}=4.2\cdot 10^{-13}$, and ${\rm maxerr}_{3}=4.9\cdot 10^{-13}$. By increasing the threshold $\delta$ to $10^{-8}$ for Algorithm 1 and $10^{-10}$ for Algorithm 3 we get $p_{1}=p_{3}=1$.
Example 6.5
In [22, Ex. 7] we find the following singular polynomial matrix of degree $8$:
$P(\lambda)={\mbox{\scriptsize$\left[\!\!\begin{array}[]{ccc}\lambda^{2}+\lambda^{8}&\lambda+\lambda^{7}&\phantom{-}\lambda^{4}\\[1.42262pt] -\lambda-\lambda^{7}&-1-\lambda^{6}&-\lambda^{3}\\[1.42262pt] \lambda^{4}&\lambda^{3}&\phantom{-}1\end{array}\!\!\right]$}},$ (18)
where $\mathop{\rm nrank}\nolimits(P(\lambda))=2$. The problem has 14 infinite eigenvalues but no finite eigenvalues. The results obtained by Algorithm 2 are presented in Table 6. The values in rows 2–11 that we left out are similar to the values in rows 1 and 12. We get two exact infinite eigenvalues in rows 13–14 because the leading matrix coefficient has rank 1, while the remaining infinite eigenvalues are numerically evaluated as simple eigenvalues $\lambda_{1},\ldots,\lambda_{12}$ with absolute values ${\cal O}(10^{1})$ that are correctly identified as infinite eigenvalues by Algorithm 4. The empirical probabilities that the methods do not return any finite eigenvalues are $p_{1}=p_{2}=p_{3}=1$ for Algorithms 1–3.
Table 6: Results of Algorithm 2 (projection to normal rank) applied to the singular PEP (18).
$j$ $\lambda_{j}$ $\quad\ \gamma_{j}$ $\quad\ \alpha_{j}$ $\quad\ \beta_{j}$ ${\rm gap}_{j}$ Type
1 $-3.01396+17.32212i$ $1.1\cdot 10^{-17}$ $1.7\cdot 10^{-16}$ $2.1\cdot 10^{-16}$ $0.51$ Infinite true
$\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$
12 $17.16068+3.09912i$ $9.0\cdot 10^{-19}$ $1.3\cdot 10^{-16}$ $2.1\cdot 10^{-16}$ $0.52$ Infinite true
13 $\infty$ $0.0$ $1.7\cdot 10^{-16}$ $6.8\cdot 10^{-16}$ $1.00$ Infinite true
14 $\infty$ $0.0$ $1.3\cdot 10^{-15}$ $2.6\cdot 10^{-16}$ $1.00$ Infinite true
15 $0.364834-0.699938i$ $1.5\cdot 10^{-1}$ $4.3\cdot 10^{-16}$ $1.2\cdot 10^{-1}$ $0.65$ Random right
16 $0.258964-1.516862i$ $2.2\cdot 10^{-3}$ $3.0\cdot 10^{-3}$ $3.3\cdot 10^{-16}$ $0.45$ Random right
7 Conclusions
We have studied singular polynomial eigenvalue problems, and extended results from singular pencils [10, 11] to this new situation.
In the PEP context, we exploit the Smith form as in Theorem 2.1 instead of the Kronecker normal form for pencils. Since there is no Kronecker canonical form for matrix polynomials under strict equivalence, we need different proof arguments. In addition, one of the main technical challenges of this paper is to show that the so-called random eigenvalues are generically simple. While there is a relatively simple and elegant argument for this in the pencil case (see [10, p. 1034]), this is much more complicated for the PEP; cf. the proof in the appendix.
As applications, we have considered certain bivariate polynomial systems, ZGV points, and several scattered examples from [15] and other references. We compared some of our results with those of [15]. An advantage of [15] is that no normal rank computation is needed; on the other hand, the methods introduced in this paper appear to be a bit more accurate. While all approaches generally perform very well, especially the projection approach, which is often also the most efficient one, seems to be slightly more reliable and accurate.
Acknowledgements
BP has been supported by the Slovenian Research and Innovation Agency (grants N1-0154, P1-0294). Parts of the results have been obtained while the first two authors visited the third author at the University of Ljubljana. MH and CM thank BP and the Faculty of Mathematics and Physics for their hospitality.
Appendix A Simple random eigenvalues
Let $P(\lambda)$ be an $n\times n$ singular matrix polynomial of degree $d$ and normal rank $n-k$, and let $S_{R}(\lambda)=\left[x_{1}(\lambda)\ \dots\ x_{k}(\lambda)\right]$ be an $n\times k$ matrix polynomial, such that its columns form a minimal basis of the right nullspace of $P(\lambda)$ with right minimal indices $m_{1},\dots,m_{k}$. In the proof of Lemma 3.1 it was already shown that generically with respect to the entries of $V^{*}\in\mathbb{C}^{k\times n}$ the polynomial $\det\big{(}V^{*}S_{R}(\lambda)\big{)}$ has exactly $M=m_{1}+\cdots+m_{k}$ roots. In this appendix we complete the proof that generically these roots are simple by providing one particular example for $V$ such that the roots of $\det\big{(}V^{*}S_{R}(\lambda)\big{)}$ are simple.
Recall that if $p_{1}(\lambda)$ and $p_{2}(\lambda)$ are two scalar polynomials, then their greatest common divisor ${\rm gcd}\big{(}p_{1}(\lambda),p_{2}(\lambda)\big{)}$ is a polynomial of maximal degree that divides both $p_{1}(\lambda)$ and $p_{2}(\lambda)$. For more than two polynomials,
${\rm gcd}\big{(}p_{1}(\lambda),\dots,p_{k}(\lambda)\big{)}={\rm gcd}\big{(}{\rm gcd}(p_{1}(\lambda),\dots,p_{k-1}(\lambda))\,,\,p_{k}(\lambda)\big{)}.$
It is well known that the greatest common divisor is unique up to multiplication with a nonzero constant.
Lemma A.1
Let $p_{1}(\lambda),\dots,p_{k}(\lambda)$ be such that their greatest common divisor $q(\lambda)={\rm gcd}\big{(}p_{1}(\lambda),\dots,p_{k}(\lambda)\big{)}$ has simple roots. Then there exists a generic set $\Omega\subseteq\mathbb{C}^{k}$ such that for all $s=(s_{1},\dots,s_{k})\in\Omega$ the roots of the polynomial $s_{1}\,p_{1}(\lambda)+\cdots+s_{k}\,p_{k}(\lambda)$ are simple.
Proof 5
Let $p_{i}(\lambda)=q(\lambda)\,h_{i}(\lambda)$ for $i=1,\ldots,k$, where ${\rm gcd}\big{(}h_{1}(\lambda),\dots,h_{k}(\lambda)\big{)}=1$. Let $\eta_{1},\dots,\eta_{\ell}\in\mathbb{C}$, where $\ell$ is the degree of $q(\lambda)$, be the roots of $q(\lambda)$, that are simple by assumption.
If $h_{1}(\lambda),\ldots,h_{k}(\lambda)$ are all constant polynomials, then $p_{1}(\lambda),\ldots,p_{k}(\lambda)$ are all multiples of $q(\lambda)$ and hence for any $s\in\mathbb{C}^{k}$ such that $s_{1}h_{1}(\lambda)+\cdots+s_{k}h_{k}(\lambda)\neq 0$ the roots of $s_{1}p_{1}(\lambda)+\cdots+s_{k}p_{k}(\lambda)$ are $\eta_{1},\dots,\eta_{\ell}$ and thus simple.
Assume now that the degree of at least one of the polynomials $h_{1}(\lambda),\ldots,h_{k}(\lambda)$ is nonzero. Suppose that $\xi$ is a multiple zero of $s_{1}h_{1}(\lambda)+\cdots+s_{k}h_{k}(\lambda)$ for some $s\in\mathbb{C}^{k}$. Then $s_{1}h_{1}(\xi)+\cdots+s_{k}h_{k}(\xi)=0$ and $s_{1}h_{1}^{\prime}(\xi)+\cdots+s_{k}h_{k}^{\prime}(\xi)=0$. It follows that $s$ is in the kernel of the $2\times k$ polynomial matrix
$H(\xi):=\left[\begin{matrix}h_{1}(\xi)&\cdots&h_{k}(\xi)\cr h_{1}^{\prime}(\xi)&\cdots&h_{k}^{\prime}(\xi)\end{matrix}\right]$ (19)
that does not depend on $s$. We know that $\big{(}h_{1}(\xi),\ldots,h_{k}(\xi)\big{)}\neq(0,\ldots,0)$ for each $\xi\in\mathbb{C}$ because ${\rm gcd}\big{(}h_{1}(\lambda),\dots,h_{k}(\lambda)\big{)}=1$, so $\mathop{\rm rank}\nolimits\big{(}H(\xi)\big{)}\geq 1$ for all $\xi$.
Let us show that $\mathop{\rm nrank}\nolimits\big{(}H(\xi)\big{)}=2$. Without loss of generality we can assume that degree of $h_{1}(\lambda)$ is at least one, so there exists $\zeta\in\mathbb{C}$ such that $h_{1}(\zeta)=0$. Since ${\rm gcd}\big{(}h_{1}(\lambda),\dots,h_{k}(\lambda)\big{)}=1$, at least one of the remaining polynomials, which we can assume to be $h_{2}(\lambda)$, is nonzero at $\zeta$. Now we consider the following polynomial, which is a minor of $H(\lambda)$,
$h_{1}(\lambda)\,h_{2}^{\prime}(\lambda)-h_{1}(\lambda)\,h_{2}^{\prime}(\lambda).$ (20)
Let $h_{i}(\lambda)=g(\lambda)r_{i}(\lambda)$ for $i=1,2$, where $g(\lambda)={\rm gcd}\big{(}h_{1}(\lambda),h_{2}(\lambda)\big{)}$ and ${\rm gcd}\big{(}r_{1}(\lambda),r_{2}(\lambda)\big{)}=1$. We know that $r_{1}(\lambda)$ is not a constant polynomial and that $r_{1}(\zeta)=0$. Suppose now that (20) is identically zero. Then
$g(\lambda)\,r_{1}(\lambda)\,\big{(}g(\lambda)\,r_{2}^{\prime}(\lambda)+g^{\prime}(\lambda)\,r_{2}(\lambda)\big{)}=g(\lambda)\,r_{2}(\lambda)\,\big{(}g(\lambda)\,r_{1}^{\prime}(\lambda)+g^{\prime}(\lambda)\,r_{1}(\lambda)\big{)}$
and it follows that $r_{1}(\lambda)\,r_{2}^{\prime}(\lambda)=r_{1}^{\prime}(\lambda)\,r_{2}(\lambda)$. But then it follows from ${\rm gcd}\big{(}r_{1}(\lambda),r_{2}(\lambda)\big{)}=1$ that $r_{1}(\lambda)$ divides its derivative $r_{1}^{\prime}(\lambda)$, which is not possible. Therefore (20) is not identically zero and, since this is a minor of $H(\lambda)$, this shows that $\mathop{\rm nrank}\nolimits\big{(}H(\lambda)\big{)}=2$.
Then there are only finitely many values $\xi_{1},\ldots,\xi_{r}$ such that $\mathop{\rm rank}\nolimits\big{(}H(\xi_{i})\big{)}=1$ for $i=1,\ldots,r$. This is how we get $r+\ell$ hyperplanes $s_{1}h_{1}(\xi_{i})+\cdots+s_{k}h_{k}(\xi_{i})=0$, $i=1,\ldots,r$, and $s_{1}h_{1}(\eta_{j})+\cdots+s_{k}h_{k}(\eta_{j})=0$, $j=1,\ldots,\ell$. If $s$ is not on any of these hyperplanes, the roots of $s_{1}h_{1}(\lambda)+\cdots+s_{k}h_{k}(\lambda)$ are simple and different from $\eta_{1},\ldots,\eta_{\ell}$. The roots of $s_{1}p_{1}(\lambda)+\cdots+s_{k}p_{k}(\lambda)$ are then simple, and since hyperplanes are algebraic sets, this holds generically for $s\in\mathbb{C}^{k}$$\qed$
Let the columns of $S_{R}(\lambda)=\left[x_{1}(\lambda)\ \dots\ x_{k}(\lambda)\right]$ form a minimal basis of the right nullspace of an $n\times n$ singular matrix polynomial $P(\lambda)$ of degree $d$ and normal rank $n-k$ with minimal indices $m_{1},\dots,m_{k}$, where we can assume that $m_{1}\leq\cdots\leq m_{k}$. Then $x_{i}(\lambda)$ is a polynomial of degree $m_{i}$ and we can write $x_{i}(\lambda)=x_{i}^{(0)}+\lambda x_{i}^{(1)}+\cdots+\lambda^{m_{i}}x_{i}^{(m_{i})}$, where $x_{i}^{(0)},\dots,x_{i}^{(m_{i})}\in\mathbb{C}^{n}$ for $i=1,\dots,k$. Then from [6] we know the following:
  1. 1)
    The high-order coefficient matrix $S_{h}=[\,x_{1}^{(m_{1})}\ \dots\ x_{k}^{(m_{k})}\,]$ has full rank.
  2. 2)
    $S_{R}(\lambda)$ has full rank for all $\lambda\in\mathbb{C}$; in particular $S_{R}(0)=[\,x_{1}^{(0)}\ \dots\ x_{k}^{(0)}\,]$ has full rank.
  3. 3)
    The greatest common divisor of all $k\times k$ minors of $S_{R}(\lambda)$ is $1$.
  4. 4)
    We have ${\rm gcd}\big{(}x_{i1}(\lambda),\dots,x_{in}(\lambda)\big{)}=1$ for all $i=1,\dots,k$.
Lemma A.2
Let $S_{R}(\lambda)=[\,x_{1}(\lambda)\ \dots\ x_{k}(\lambda)\,]$ be an $n\times k$ matrix polynomial, such that its columns form a minimal basis of the right nullspace of $P(\lambda)$ with right minimal indices $m_{1},\dots,m_{k}$, where $k\geq 2$. Then, there exists vectors $v_{1},\dots,v_{k-1}\in\mathbb{C}^{n}$ such that
$\mathop{\rm rank}\nolimits\big{(}[v_{1}\,\dots\,v_{k-1}]^{*}S_{R}(\lambda)\big{)}=k-1\ \textrm{for all }\ \lambda\in\mathbb{C}.$ (21)
Proof 6
We apply induction and start with $k=2$. Thus, we aim to find $v_{1}\in\mathbb{C}^{n}$ such that $\operatorname{rank}\big{(}[v_{1}^{*}x_{1}(\lambda)\ \ v_{1}^{*}x_{2}(\lambda)]\big{)}=1$ for all $\lambda\in\mathbb{C}$. Generically with respect to the entries of $v_{1}$ the polynomials $v_{1}^{*}x_{1}(\lambda)$ and $v_{1}^{*}x_{2}(\lambda)$ will have $m_{1}$ and $m_{2}$ roots, respectively. Let us select such a $v_{1}\in\mathbb{C}^{n}$ and let $\alpha_{1},\dots,\alpha_{m_{1}}$ denote the roots of $v_{1}^{*}x_{1}(\lambda)$ and $\beta_{1},\dots,\beta_{m_{2}}$ those of $v_{1}^{*}x_{2}(\lambda)$. (We highlight that enforcing the maximal possible degree of each polynomial $v_{i}^{*}x_{j}(\lambda)$ ensures that the number of roots remains invariant under sufficiently small perturbations.) Suppose that the roots are ordered in such a way that
$\alpha_{i}=\beta_{i}\quad\mbox{for }i=1,\dots,r\quad\mbox{ and }\quad\alpha_{i}\neq\beta_{i}\quad\mbox{for }i=r+1,\dots,\min(m_{1},m_{2}).$
In other words, we have exactly $r$ solutions $\lambda_{i}:=\alpha_{i}=\beta_{i}$, $i=1,\dots,r$, to the equation
$v_{1}^{*}x_{1}(\lambda)=v_{1}^{*}x_{2}(\lambda)=0.$ (22)
[…]
Assume now that the statement holds for $k-1$. Then for $\widehat{S}_{R}(\lambda)=[\,x_{1}(\lambda)\ \dots\ x_{k-1}(\lambda)\,]$ there exist $k-2$ vectors $v_{1},\dots,v_{k-2}\in\mathbb{C}^{n}$ such that
$\operatorname{rank}\big{(}[v_{1}\ \dots\ v_{k-2}]^{*}\widehat{S}_{R}(\lambda)\big{)}=k-2$ (23)
for all $\lambda\in\mathbb{C}$. Without loss of generality we may assume that each $v_{i}^{*}x_{j}(\lambda)$ has degree $m_{j}$ as a polynomial in $\lambda$. Note that this is a generic condition, so if, by chance, this is not satisfied, then a small perturbation of the vectors $v_{1},\dots,v_{k-2}$ will ensure that the condition is satisfied. If the perturbation is small enough then this does not change the rank condition (23). Next, select $v_{k-1}\in\mathbb{C}^{n}$ such that $v_{k-1}^{*}x_{j}(\lambda)$ has $m_{j}$ zeros for $j=1,\dots,k$. (Note that the latter condition is again generic, so such an $v_{k-1}$ exists.) Then by an argument similar to the one in the case $k=2$ there can be only finitely many distinct values $\lambda_{1},\dots,\lambda_{r}$ that are solutions of
$\mathop{\rm rank}\nolimits\big{(}[v_{1}\ \dots\ v_{k-1}]^{*}S_{R}(\lambda)\big{)}=k-2.$ (24)
Indeed, we have that the rank in (24) is less then $k-1$ if all $(k-1)\times(k-1)$ minors are zero and there is only a finite number of common zeros of these polynomials. Also note that the rank in (24) cannot be smaller than $k-2$, because of (23). In a same way as for $k=2$, if $r=0$, we are done. Otherwise, since the vectors $x_{1}(\lambda_{1}),\dots,x_{k}(\lambda_{1})$ are linearly independent, there exists vector $w_{k-1}$ such that $w_{k-1}^{*}x_{j}(\lambda_{1})=0$ for $j=1,\dots,k-2$ and $w_{k-1}^{*}x_{k-1}(\lambda_{1})\neq 0$. For a sufficiently small $\delta_{k-1}>0$ we then have
$[v_{1}\ \dots\ v_{k-2}\ \ v_{k-1}+\delta_{k-1}w_{k-1}]^{*}S_{R}(\lambda_{1})$ $=[v_{1}\ \dots\ v_{k-1}]^{*}S_{R}(\lambda_{1})$
$\phantom{MMM}+\delta_{k-1}\,w_{k-1}^{*}x_{k-1}(\lambda_{1})\,e_{k-1}\,e_{k}^{T}$ (25)
and the rank of (25) is $k-1$. If we replace $v_{k-1}$ by $v_{k-1}+\delta_{k-1}w_{k-1}$ then $\lambda_{1}$ is not a solution of (24) anymore and due to the continuity of zeros, there can now be at most $r-1$ values $\lambda$ such that (24) holds. We can remove them in the same way one by one by further perturbing $v_{k-1}$. This delivers the desired example for which (21) holds.$\qed$
Theorem A.3
Let $P(\lambda)$ be an $n\times n$ singular matrix polynomial of degree $d$ and normal rank $n-k$, and let $S_{R}(\lambda)=\left[x_{1}(\lambda)\ \dots\ x_{k}(\lambda)\right]$ be an $n\times k$ matrix polynomial, such that its columns form a minimal basis of the right nullspace of $P(\lambda)$ with right minimal indices $m_{1},\dots,m_{k}$. Then, there exists $V\in\mathbb{C}^{n\times k}$ such that the polynomial $\det\big{(}V^{*}S_{R}(\lambda)\big{)}$ has exactly $M=m_{1}+\cdots+m_{k}$ simple zeros.
Proof 7
For $k=1$ we can take $V=v\in\mathbb{C}^{n}$ and let $g(\lambda):=V^{*}S_{R}(\lambda)=\sum_{j=1}^{m_{1}}{v_{j}^{*}x_{1j}(\lambda)}$. Since ${\rm gcd}\big{(}x_{11}(\lambda),\dots,x_{1m_{1}}(\lambda)\big{)}=1$, it follows from Lemma A.1 that generically the polynomial $g(\lambda)$ has exactly $m_{1}$ simple zeros.
For $k\geq 2$, select $v_{1},\dots,v_{k-1}\in\mathbb{C}^{n}$ satisfying (21) according to Lemma A.2 and an additional $v_{k}\in\mathbb{C}^{n}$, and set $G_{1}(\lambda)=[v_{1}\ \dots\ v_{k}]^{*}S_{R}(\lambda)$. Without loss of generality we may assume that $g_{1}(\lambda)=\det(G_{1}(\lambda))$ has degree $M=m_{1}+\cdots+m_{k}$. Indeed, since this is a generic condition, we can slightly perturb $v_{1},\dots,v_{k}$ to meet the condition, in the case that, by chance, this is not satisfied. If the perturbation is sufficiently small, then this will not change the rank condition (21). It follows that $g_{1}$ has exactly $M$ roots $\xi_{1},\dots,\xi_{M}$. At this point, we do not know if the roots are pairwise distinct. Therefore, we next consider $G_{2}(\lambda)=[v_{1}\ \dots\ v_{k-1}\ \ w]^{*}S_{R}(\lambda)$, where the vectors $v_{1},\dots,v_{k-1}$ are as before and $w\in\mathbb{C}^{n}$. Again, we can assume without loss of generality that $g_{2}(\lambda)=\det G_{2}(\lambda)$ has degree $M$. By construction, the first $k-1$ rows of $G_{2}(\lambda)$ have full rank for all $\lambda$. Consequently, a fixed value $\xi$ will generically not be among the roots of $g_{2}(\lambda)=\det(G_{2}(\lambda))$ as a polynomial in $\lambda$ for $i=1,\dots,M$. Indeed, since $S_{R}(\xi)$ has full rank $k$, the range of $S_{R}(\xi)^{*}$ is $k$-dimensional and $S_{R}(\xi)^{*}v_{1},\dots,S_{R}(\xi)^{*}v_{k-1}$ are linearly independent. Choosing $w$ such that $S_{R}(\xi)^{*}w$ completes these vectors to a basis of the range of $S_{R}(\xi)^{*}$ proves that $G_{2}(\xi)$ is nonsingular. But then the set of all $w$ for which $g_{2}(\xi)\neq 0$ is a generic set. Since the intersection of finitely many generic sets is still generic, it follows that we can generically exclude all values $\xi_{1},\dots,\xi_{M}$ from the zeros $\eta_{1},\dots,\eta_{M}$ of $g_{2}(\lambda)$.
It follows that ${\rm gcd}\big{(}g_{1}(\lambda),g_{2}(\lambda)\big{)}=1$ and we can apply Lemma A.1. Therefore, there exists $(s_{1},s_{2})\in\mathbb{C}^{2}$ such that $s_{1}g_{1}(\lambda)+s_{2}g_{2}(\lambda)$ has $M$ simple zeros. If we take $V=[v_{1}\ \dots\ v_{k-1}\ \ \overline{s_{1}}v_{k}+\overline{s_{2}}w]$, then $\det\big{(}V^{*}R_{S}(\lambda)\big{)}$ has $M$ simple roots which completes the proof.$\qed$